### v1

# Meta-learning of Gibbs states for many-body Hamiltonians with applications to Quantum Boltzmann Machines

Abstract
The preparation of quantum Gibbs states is a fundamental challenge in quantum computing, essential for applications ranging from modeling open quantum systems to quantum machine learning. Building on the Meta-Variational Quantum Eigensolver framework proposed by Cervera-Lierta et al.[1] and a problem-driven ansatz design, we introduce two meta-learning algorithms : Meta-Variational Quantum Thermalizer (Meta-VQT) and Neural Network Meta-VQT (NN-Meta VQT) for efficient thermal state preparation of parametrized Hamiltonians on Noisy Intermediate-Scale Quantum (NISQ) devices. Meta-VQT utilizes a fully quantum ansatz, while NN-Meta VQT integrates a quantum-classical hybrid architecture. Both leverage collective optimization over training sets to generalize Gibbs state preparation to unseen parameters. We validate our methods on upto 8-qubit Transverse Field Ising Model and the 2-qubit Heisenberg model with all field terms, demonstrating efficient thermal state generation beyond training data. For larger systems, we show that our meta-learned parameters when combined with appropriately designed ansatz serve as warm-start initializations, significantly outperforming random initializations in the optimization tasks. Furthermore, a 3-qubit Kitaev ring example showcases our algorithm’s effectiveness across finite-temperature crossover regimes. Finally, we apply our algorithms to train a Quantum Boltzmann Machine (QBM) on a 2-qubit Heisenberg model with all field terms, achieving enhanced training efficiency, improved Gibbs state accuracy, and a 30-fold runtime speedup over existing techniques such as variational quantum imaginary time (VarQITE)-based QBM highlighting the scalability and practicality of meta-algorithm-based QBMs.
I Introduction
Generation of finite-temperature thermal states or Gibbs states of a given Hamiltonian is an integral task in state preparation using quantum computers. Gibbs states find applications in a variety of tasks such as quantum simulation [2], quantum machine learning [3, 4], quantum optimization [5], and the study of open quantum systems [6]. In particular, combinatorial optimization problems [5], semi-definite programming [7], quantum key distribution [8], quantum metrology [9], and training of Quantum Boltzmann Machines (QBMs) [3, 10, 11, 12, 13, 14, 15] are tackled by sampling from well-engineered Gibbs states.
Preparing arbitrary quantum states is generally a challenging task; in particular, finding the ground state of a Hamiltonian is known to be QMA-hard [16]. Preparing Gibbs states, especially at low and intermediate temperatures, is as hard as finding the ground state of that Hamiltonian [17]. Various methods propose preparing Gibbs states using classical and quantum computers [18, 6, 19, 20, 21, 22, 23, 24]. Some proposals that focus on quantum computers include quantum rejection sampling [25], quantum walk [26], and dynamics simulation [27, 28]. Although these methods may incur exponential cost in the worst case, they are efficient in special scenarios such as when the ratio between the partition functions of the infinite temperature state and the Gibbs state is at most polynomially large [6], and when the gap of the Markov chain describing the quantum walk is polynomially small [27, 26]. However, these algorithms suit fault-tolerant quantum computers as they require complex quantum subroutines like quantum phase estimation.
Figure 1: Overview of meta-algorithm for Gibbs state preparation :(a) Depicts the core task addressed by our meta-algorithm: preparing thermal states for a parametrized many-body spin Hamiltonian at finite temperature. Shown are illustrative profiles of the Gibbs free energy and the purity of the corresponding Gibbs state as functions of a Hamiltonian parameter. The parameter space is split into training (green circles) and test (blue star) sets. (b) Illustrates the training process, where parameters from the training set are encoded into a variational ansatz composed of a hardware-efficient and a Hamiltonian variational ansatz, separated by a linear arrangement of CNOT gates. The global cost, evaluated across all training points, is minimized using a classical optimizer. (c) Shows inference, where test parameters are input into the trained ansatz, initialized with the optimized parameters from (b), to generate Gibbs states. These are then used to reconstruct the Gibbs free energy profile.
Variational Quantum Algorithms (VQAs) that can be implemented efficiently on noisy intermediate-scale quantum (NISQ) devices, present an alternative to these methods. VQAs are a class of hybrid quantum-classical algorithms that optimize a loss function depending on the circuit parameters. The loss function is evaluated on quantum hardware, while parameter updates are computed on classical processors.
Several variational techniques have also been proposed for Gibbs state preparation on near-term quantum hardware. These include Variational Quantum Thermalizers (VQTs) which are the finite-temperature extensions of the variational quantum eigensolver (VQE) [29, 30, 31, 32], methods that utilize variational quantum imaginary time evolution (VarQITE) [33, 34, 11], and truncated Taylor series expansions of the Gibbs free energy as a cost function [35]. Additionally, dynamic parameterized quantum circuits (PQCs) that mitigate barren plateaus [36], and non-unitary multi-qubit operations [37] are proposed to enhance algorithm performance for Gibbs state preparation. A significant drawback of these algorithms is that they must be executed separately for each Hamiltonian instance, making them computationally expensive when preparing Gibbs states for a family of parameterized Hamiltonians—for example, during the training of QBMs [10, 11], or when exploring thermal properties across varying coupling strengths to construct many-body phase diagrams [38, 39]. Another challenge is to select an ansatz with sufficient expressivity to efficiently prepare the Gibbs state for specific Hamiltonian parameters and temperatures, especially when the Hamiltonian undergoes finite-temperature crossovers as a function of those parameters [40].
Table 1: System and ancilla qubit counts in prior variational approaches to Gibbs state preparation.
Prior Variational Approaches
System
Qubits
Ancilla
Qubits
Quantum Hamiltonian-Based Models and the Variational Quantum Thermalizer Algorithm, Verdon et al. [29] $4$ 0
Variational quantum Gibbs state preparation with a truncated Taylor series, Wang et al.[35] 11 1
Extending the Variational Quantum Eigensolver to Finite Temperatures, Selisko et al.[31] 4 0
Variational quantum simulation of the quantum critical regime, Shi et al. [40] 6 0
Training Quantum Boltzmann Machines with the $\beta$ Variational Quantum Eigensolver, Huijgen et al.[41] 10 0
Thermal state preparation of the SYK model using a variational quantum algorithm, Araz et al.[32] 12 0
Dynamic parameterized quantum circuits: expressive and barren-plateau free, Deshpande et al.[36] 10 0
Meta and Neural Network Meta-Variational Quantum Thermalizer, Our Work 8 8
In this work, we address these challenges by proposing two novel algorithms :
  • Meta-Variational Quantum Thermalizer (Meta-VQT) and
  • Neural Network Meta-VQT (NN Meta-VQT),
for preparing the quantum Gibbs state of a parameterized Hamiltonian. While inspired by Cervera-Lierta et al.[1], Miao et al. [42], and related collective optimization approaches such as Zhang and Yin [43] and Mitarai et al. [44], which focus on learning ground state (zero temperature) energy profiles, our algorithms are distinct and specifically tailored for finite-temperature Gibbs state preparation. The ansatz choice in both Meta-VQT and NN Meta-VQT algorithms captures the complex correlations appearing in the Gibbs state. Both algorithms train using a collective optimization (meta-learning) technique where a finite set of Hamiltonian parameters are encoded into the ansatz along with trainable angles. A global cost function is constructed from individual cost functions defined for each Hamiltonian parameter. Optimizing this cost function allows the ansatz, after training, to approximate the quantum Gibbs state of the Hamiltonian for parameters beyond the training set. The schematic workflow of our task, meta-learning, and generation of the Gibbs state and Gibbs free energy profile for unseen points from the trained ansatz is shown in Fig. 1.
Figure 2: Workflow of VQT, Meta-VQT and NN Meta-VQT : The workflow of VQT, where the PQC is trained for approximating the Gibbs state corresponding to a particular parameter value, $\vec{h}_{1}$, of the Hamiltonian is shown in (a). For a different parameter, this workflow needs to be run again. The workflow of Meta-VQT and NN Meta-VQT is shown in (b). Unlike (a) here a set of Hamiltonian parameters $\vec{h}_{train}$ (Step 1), are encoded in the PQC (Step 3, pink colored box) or classical neural network (Step 3, green colored box), and a global loss function is evaluated (Step 7) that does collective optimization over these set of parameters. After training, the PQC can take any value in $\vec{h}_{test}\notin\vec{h}_{train}$ as well as those that lies within the range of $\vec{h}_{train}$ as input and accurately prepare Gibbs state corresponding to these parameters without the need for further training. In both (a) and (b), the blue and orange boxes represent modules that are run on classical and quantum computers respectively.
To demonstrate the efficacy of our proposed meta-algorithms, we investigate their performance across a diverse set of parameterized quantum many-body systems. Specifically, we consider the following representative models:
  • a single parametrized Transverse Field Ising Model (TFIM) of sizes from $n=2$ to $8$ qubits,
  • single parametrized Kitaev ring model of size $n=3$ qubits, that is known to exhibit finite-temperature crossover for smaller system sizes [45],
  • double parametrized $2$-qubit Heisenberg Hamiltonian with all field terms that has $3$ commuting blocks.
In all cases, we use an equal number of ancilla and system qubits. The total number of qubits involved is therefore comparable to the scales explored in existing variational Gibbs state preparation studies (see Table 1). We provide numerical evidence that $(i)$ after training, our algorithm can prepare quantum states that closely approximate the true Gibbs states of parametrized Hamiltonians even for parameter values not included in the training set, and (ii) if the estimation is not precise enough then the meta-trained circuit can be used as the starting point of a traditional VQT rather than random initializations.
Finally, we demonstrate the utility of these algorithms by improving the training of a variational QBM with a simple example: a 2-qubit double-parameterized Heisenberg Hamiltonian with external fields. Upon comparison of our results with an existing technique of VarQITE-based variational QBM, we observe that our collective optimization algorithms offer several advantages. Specifically,
  • Enhance the training speed and reduce the complexity by requiring fewer QPU calls,
  • Increase the training efficiency by more accurate Gibbs state preparation, as the system size increases and the Hamiltonian and input data distributions become more complex.
For relatively simple Hamiltonians with fewer non-commuting terms, our approach has performance comparable to existing methods. However, as the complexity of the Hamiltonian grows, our proposed protocol have the potential to offer scalable and efficient performance, highlighting its broader applicability and robustness for practical quantum machine learning tasks.
Figure 3: Schematic workflow of meta-learning algorithms : In the workflow presented here, different types of training data denoted by training data 1,..,training data N are associated with distinct learning algorithms and models, each defined by a separate loss function. The meta-learning approach optimizes a combined loss function based on $P_{i,j}$ and shared feature representations and configurations, $m_{j}~{}\text{and}~{}\lambda_{i}$, enabling the resulting meta-learner model to perform effectively on unseen new data that was not included in the training dataset. This schematic workflow has been adapted from Bahri et al.[46].
II Background
In this section, we briefly review one of the most widely studied NISQ compatible approaches for the preparation of Gibbs states: the VQT introduced by Verdon et al.[29], which forms the foundation of our proposed meta-algorithms. The VarQITE framework developed by Zoufal et al. [11], which we use as a baseline for comparative analysis throughout this work is explained briefly in Appendix A.
II.1 Variational Quantum Thermalizer (VQT)
Consider a n-qubit Hamiltonian $H(\vec{h})$ at temperature $T$. Then the Gibbs state of this parametrized Hamiltonian is given as,
$\rho(\vec{h})=\frac{e^{-\beta H(\vec{h})}}{\sum e^{-\beta H(\vec{h})}},$ (1)
where $\beta=1/k_{B}T$ with $T$ being the temperature, $k_{B}$ is the Boltzmann constant and $\sum e^{-\beta H(\vec{h})}$ is the partition function. For convenience, we assume that the Boltzmann constant is equal to $1$ throughout the paper.
Given a PQC represented by unitaries $U(\vec{\theta})$ with $m$ trainable angles, $\vec{\theta}\in\theta_{1},\theta_{2},...\theta_{m}$ acting on $\text{n}+\text{n}_{\text{ancilla}}$ qubits, the goal of VQT is to train the PQC to approximate the Gibbs state of this Hamiltonian as, $\rho(\vec{h})\approx\tilde{\rho}(\vec{\theta})=\text{Tr}_{\text{ancilla}}\left(\ket{\psi(\vec{\theta})}\bra{\psi(\vec{\theta})}\right)$, where |$\ket{\psi(\vec{\theta})}=U(\vec{\theta})\ket{\psi_{0}}$ is the state generated by the PQC from the initial state |$\ket{\psi_{0}}$. The exact loss function to be minimized using a classical computer to train this PQC is the Gibbs free energy given by,
$G(\vec{\theta},\vec{h})=\langle H(\vec{h})\rangle-TS,$ (2)
where $\langle H(\vec{h})\rangle=\text{Tr}[H(\vec{h})\tilde{\rho}(\vec{\theta})]$ is the expectation value of the Hamiltonian, $S=-\text{Tr}[\tilde{\rho}(\vec{\theta})\log\tilde{\rho}(\vec{\theta})]$ is the Von-Neumann entropy and $T=1/\beta$ is the temperature of the system. The minimization of $G(\vec{\theta},\vec{h})$ ensures that the system reaches the corresponding Gibbs state which is the state of minimal energy at finite-temperature. This is in analogy with the VQE where the expectation value of the Hamiltonian is minimized to find the ground state energy. In the limit $T\rightarrow 0$, Eq. (2) reduces to minimization of the expectation value of the Hamiltonian, which is the loss function in VQE.
However, unlike VQE, the mixed nature of the Gibbs state necessitates the use of ancilla qubits $(1~{}\leq~{}\text{n}_{\text{ancilla}}~{}\leq~{}\text{n})$ to prepare the approximate Gibbs state of given Hamiltonian. The required ancilla qubits can scale up further depending on how the loss function is evaluated on a quantum computer. For instance, Wang et al.[35] employs a truncated Taylor series expansion of order $3$ that needs a SWAP test to measure the approximate Von-Neumann entropy, thus requiring multiple copies of the quantum state and consequentially more ancillas. The schematic workflow of VQT is shown in Fig. 2(a).
Figure 4: Meta-VQT ansatz : The Hamiltonian parameters, $\vec{h}$, are encoded in the angles of the single qubit parametrized gates through a linear transformation. This along with the linear implementation of CNOTs forms the encoding layer. The output state from the encoding layer is further processed by the parametrized HVA. The gates in HVA are chosen according to the Hamiltonian whose Gibbs states needs to be prepared. An example HVA for a transverse field Ising model is shown, where $XX=e^{-i\theta\sigma_{i}^{x}\sigma_{i+1}^{x}}$. $l$ represents the number of encoding or processing layers. After the final processing layer, measurement on the system qubits yield a density matrix, $\tilde{\rho}(\vec{\omega},\vec{\phi},\vec{\theta})$, that is further optimized to approximate the true Gibbs state of the system. The number of ancillas are taken to be same as the number of system qubits.
III Methods
In this section, we introduce the finite-temperature collective optimization using two meta-algorithms developed for Gibbs state preparation: the Meta-VQT and the NN Meta-VQT.
III.1 Finite-temperature collective optimization
In classical machine learning, collective optimization or meta-learning (learning to learn) involves training an outer (or meta) algorithm that updates the inner learning algorithm. This inner algorithm, used for tasks like image classification defined by a dataset and objective, is refined by the meta-algorithm to enhance the model’s performance on broader objectives, such as improving generalization or accelerating the learning speed of the inner algorithm. Meta-learning also draws on prior experience to handle new tasks more effectively [47, 48, 49, 46]. A schematic diagram that summarizes the meta-learning algorithm is shown in Fig. 3.
Recently, meta-learning has attracted significant attention from the quantum machine learning community as well. It finds applications in tasks such as generalization of ground states of parametrized Hamiltonian (Meta-VQE and Neural Network (NN) Meta-VQE or quantum circuit learning) [1, 42, 43, 44], training algorithms that does not rely on gradient computation [50] and training classical neural networks to assist in the quantum learning process [51, 52].
In this work, we introduce the finite-temperature versions of Meta-VQE and NN Meta-VQE, known as Meta-VQT and NN Meta-VQT, which generates the Gibbs free energy profile, $G(\vec{h})$, and prepares the corresponding quantum Gibbs state, $\rho(\vec{h})$, of a parametrized Hamiltonian, $H(\vec{h})$. The workflow of our proposed Meta-VQT/NN Meta-VQT is shown in Fig. 2(b). The finite-temperature meta-algorithms are developed as generalized versions of the VQT algorithm explained in Sec. II.1. Here, the PQC, |$\ket{\psi(\vec{\theta},\vec{\phi},\vec{\omega})}$ (or neural network) is trained for multiple discrete input parameters ${\vec{h}_{train}}\in[\vec{h}_{1},\vec{h}_{2},...,\vec{h}_{n}]$ of the Hamiltonian, $H(\vec{h})$, so that for the optimal ansatz parameters, $(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})$ (or $\vec{\phi}_{nn}^{*})$, the quantum circuit is able to generate quantum Gibbs state for Hamiltonian parameters $\vec{h}_{test}\notin{\vec{h}_{train}}$. Suppose that $\vec{h}$ is a single-valued vector $h$ and $h_{train}\in(h_{1},h_{2},.....h_{n})$ is the training set with $h$ values arranged in ascending order. Then the meta-algorithms perform well for $h_{test}\notin h_{train}$ and $h_{1}\leq h_{test}\leq h_{n}$, whereas outside this range, its performance tends to decline. The number of distinctive trainable angles depends on the type of meta-algorithm, size of $\vec{h}$, the type of encoding and the class of PQC used for training.
Since, the target Gibbs state is usually mixed, additional ancilla qubits are required to prepare quantum Gibbs state that we trace out before measurement. Having equal number of ancilla and system qubits provides the best performance. However, for simpler Hamiltonians, fewer ancilla qubits also demonstrates high accuracy.
Figure 5: NN Meta-VQT ansatz : In NN Meta-VQT, the neural network encodes the Hamiltonian parameters, $\vec{h}$, and provides the circuit parameters, $\vec{\theta}$, as output. The quantum circuit is constructed out of both hardware-efficient and HVA as shown above. Similar to Meta-VQT, after this final processing layer, measurement on the system qubits yield a density matrix, $\tilde{\rho}(\vec{\theta})$, that is further optimized to approximate the true Gibbs state of the system.
The loss function used is the sum of Gibbs free energy corresponding to each parameter in $\vec{h}_{train}$. For system sizes considered here, where the number of qubits ranges from $n=2~{}\text{to}~{}8$, the Gibbs free energy is evaluated on a classical computer after measuring the Gibbs state, $\rho(\vec{h})\approx\tilde{\rho}(\vec{\theta})=\text{Tr}_{\text{ancilla}}\left(\ket{\psi(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})}\bra{\psi(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})}\right)$, from the quantum circuit. For larger system sizes, where the size of the Gibbs state becomes too large to be stored in a classical computer, the Gibbs free energy itself is evaluated from the quantum circuits using a truncated Taylor series expansion of the entropy [35] or by applying the method described by Verdon et al. [29].
The following sections outline the details regarding the Meta-VQT and NN Meta-VQT algorithms.
III.1.1 Meta-Variational Quantum Thermalizer (Meta-VQT)
For the parametrized spin Hamiltonian, $H(\vec{h})$, where $\vec{h}=(h^{1},h^{2},\ldots h^{p})$ represents a vector comprising $p$ parameters of the Hamiltonian that define the interaction strength and/or field strength, consider a discrete finite set of parameters defined by $\vec{h}_{\text{train}}\in[\vec{h}_{1},\vec{h}_{2},\ldots,\vec{h}_{m}]$. The first step in the Meta-VQT involves encoding the parameters of the Hamiltonian in $\vec{h}_{train}$ in the quantum circuit, see pink colored box Fig. 2(b). This encoding layer is chosen to be a SU2 hardware efficient ansatz. This ansatz is constructed using native single-qubit and two-qubit gates, typically following the SU(2) group operations, and is optimized to align with the physical connectivity of the quantum processor [53]. The number of ancillas in the encoding layer is considered same as that of the system qubits. The encoding is done using a linear transformation in the angles of the RZ and RY quantum gates, as shown in Fig. 4, followed by a linear implementation of CNOTs. The variables $\vec{\omega}$ and $\vec{\phi}$ are the trainable parameters in the encoding layer. If $l$ is the number of encoding layers, then the total number of trainable parameters in the encoding layer is $4\times(\text{n}+\text{n}_{\text{ancilla}})\times l\times\text{size of}~{}\vec{h}$.
It is proven in several works that the SU2 hardware-efficient ansatz alone in VQT remains insufficient for simulating the quantum Gibbs state of complex Hamiltonians, especially near the quantum critical regime [40]. Several ansatze such as the product spectrum ansatz [54] and the Hamiltonian variational ansatz (HVA) [40, 29, 30, 31, 32] have been proposed to efficiently simulate the quantum Gibbs state of these complex many-body Hamiltonians. With this motivation, the quantum state from the encoding layer is further processed using HVA with trainable parameters followed by a linear implementation of CNOTs, as shown in Fig. 4. The total number of qubits in the HVA layer is also taken as $\text{n}+\text{n}_{\text{ancilla}}$. The HVA ansatz, $U(\vec{\theta})=e^{-i\vec{\theta}H(\vec{h})}$ have parametrized $1$ and $2$ qubit quantum gates chosen according to the Hamiltonian terms under consideration. The example of HVA for TFIM is shown in Fig. 4. The number of trainable parameters, $\vec{\theta}$, in HVA depends on the terms in the Hamiltonian whose Gibbs state needs to be prepared.
Figure 6: Gibbs state preparation of TFIM using Meta-VQT : The results of Meta-VQT algorithm run for different qubit sizes of TFIM is shown. The value of $\beta$ is chosen to be $1$. (a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set. The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of Meta-VQT is shown in (b). The relative error for all qubit sizes are low indicating the high efficiency of trained Meta-VQT in predicting the Gibbs free energy for test parameters outside the training set. This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from the trained Meta-VQT for test parameters gives value between $0.93$ to $1$. The number of encoding and processing layers chosen for system sizes $n=2,3,4$ are $[2,2],[3,3]~{}\text{and}~{}[4,4]$ respectively.
The density matrix, $\tilde{\rho}(\vec{\omega},\vec{\phi},\vec{\theta},\vec{h})$, measured by tracing out the ancillas is used to evaluate the Gibbs free energy, $G(\vec{h})$, of $H(\vec{h})$. The trainable parameters, $\vec{\omega},\vec{\phi},\vec{\theta}$, in the encoding and processing layers are initialized randomly at the first iteration. With the same random initialization, all the parameters in $\vec{h}_{\text{train}}$ are encoded step by step and corresponding $\tilde{\rho}(\vec{\omega},\vec{\phi},\vec{\theta},\vec{h})$ and $G(\vec{h})$ are evaluated, see Fig. 2(b). The global loss function, $G_{\text{total}}$, is then evaluated as,
$G_{\text{total}}=\sum_{\vec{h}_{i}\in\vec{h}_{\text{train}}}G(\vec{h}_{i}),\\ =\sum_{\vec{h}_{i}\in\vec{h}_{\text{train}}}\text{Tr}[H(\vec{h}_{i})\tilde{\rho}(\vec{\omega},\vec{\phi},\vec{\theta},\vec{h}_{i})]+\\ \frac{1}{\beta}\text{Tr}[\tilde{\rho}(\vec{\omega},\vec{\phi},\vec{\theta},\vec{h}_{i})\log\tilde{\rho}(\vec{\omega},\vec{\phi},\vec{\theta},\vec{h}_{i})].$ (3)
The $G_{\text{total}}$ is minimized, and the trainable parameters are updated using the classical optimizer ADAM until convergence is reached. Once converged, the quantum state measured at optimum parameters, $\vec{\omega}^{*},\vec{\phi}^{*},\vec{\theta}^{*}$, of the quantum circuit, by tracing out ancillas, is expected to be close to the true quantum Gibbs state of the given Hamiltonian at parameters $\vec{h}_{j}\in\vec{h}_{\text{train}}$ as well as $\vec{h}_{j}\notin\vec{h}_{\text{train}}$.
The chosen ansatz plays a crucial role in accurately representing thermal correlations in many-body systems. The Gibbs state of many-body quantum systems can display growing mutual information with subsystem size, driven by both the Hamiltonian parameters and temperature, and particularly pronounced near quantum critical points. Through numerical simulations we observe that our composite ansatz, which interleaves hardware-efficient (HEA) and Hamiltonian variational (HVA) layers with CNOT gates, successfully captures this behavior with high fidelity. This hybrid structure enables faithful approximation of complex Gibbs states across diverse values of $\vec{h}$ and $\beta$.
III.1.2 Neural Network Meta-Variational Quantum Thermalizer (NN Meta-VQT)
Neural Network Meta-VQT (NN Meta-VQT) builds on the previous variant by adding a classical neural network to the quantum circuit as shown in Step 3 (green colored box) of Fig. 2(b). The neural network receives the Hamiltonian parameters from $\vec{h}_{train}$, as input and provides the parameters of the quantum circuit, $\vec{\theta}$, as output, essentially taking the place of the encoding layer. This quantum circuit that takes the parameters from neural network forms the processing layer that gives the approximation to Gibbs state of the Hamiltonian. Unlike Meta-VQT, here the trainable parameters, $\vec{\phi}_{nn}$, are present in the neural network rather than in the quantum circuit. The number of trainable parameters (size of $\vec{\phi}_{nn}$) depends on the type of neural network, number of hidden layers considered and the type of processing layers included in the ansatz. Also, the number of classical trainable parameters are greater in number than the trainable parameters in Meta-VQT, thus providing an additional layer of complexity in mapping Hamiltonian parameters to the quantum circuit [42]. At the end of training, the neural network optimizes the mapping between the Hamiltonian parameters $H(\vec{h})$ and the necessary parameters of the quantum circuit required to generate the Gibbs state.
The choice of processing layer considered here includes SU2 hardware-efficient ansatz followed by HVA, as shown in Fig. 5. The neural network architecture considered here includes $3$ fully-connected hidden layers with sigmoid activation function and a linear layer at the output that matches the dimensions of the parameterized quantum circuit. During training, the global loss function, $G_{\text{total}}$ in Eq. (3), which is obtained by summing over Gibbs free energy corresponding to each of the parameters in the training set, is minimized until convergence.
Figure 7: Gibbs state preparation of TFIM using NN Meta-VQT : The results of NN Meta-VQT algorithm for different qubit sizes of TFIM is shown. The value of $\beta$ is chosen to be $1$. (a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set. The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of NN Meta-VQT is shown in (b). The relative error for all qubit sizes are low indicating the high efficiency of trained neural network architecture in predicting the Gibbs free energy for test parameters outside the training set. This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from trained NN Meta-VQT ansatz for test parameters gives value between $0.98$ to $1$ for $n=2,3$ and $4$. The number of SU2 hardware efficient and HVA ansatz layers in the processing part is chosen as $[2,2],~{}[3,3]$ and $[4,4]$ for system sizes $n=2,3$ and $4$ respectively.
Figure 8: Scaling of Meta-VQT and NN Meta-VQT:(a) shows the mean value with standard deviation of relative error between the Gibbs free energy evaluated for values of $h$ in $h_{test}$ from the trained meta-algorithms and the exact one for system sizes $2,4,6$ and $8$ qubits. The relative error increases for both Meta-VQT and NN Meta-VQT as system size increases to $8$, with NN Meta-VQT exhibiting lower value than Meta-VQT. (b) shows the mean value and standard deviation of fidelity evaluated between prepared and true Gibbs state at $h_{test}$ as a function of system size. The fidelity decreases for both Meta-VQT and NN Meta-VQT as system size increases to $8$, with a slighltly higher value for NN Meta-VQT than Meta-VQT for larger system sizes. For $n$ qubit Hamiltonian, the number of encoding (SU2 hardware efficient) and processing (HVA) layers considered for Meta-VQT (NN Meta-VQT) is $[n,n]$. The test set, $h_{test}$, over which the mean and standard deviation in relative error in Gibbs free energy and fidelity is evaluated are $40$ uniform points between $[-2,2]$.
IV Numerical Results
This section provides results from the numerical experiments where the Meta-VQT and NN Meta-VQT algorithms are used for preparing Gibbs state of different types of spin Hamiltonians. The pseudocode for deploying the Meta and NN Meta-VQT is shown under Algorithm 1 and Algorithm 2 in Appendix E.
IV.1 Preparing Gibbs state of TFIM
In this section, Meta and NN Meta-VQT algorithms are deployed to generate the quantum Gibbs state of a parametrized transverse field ising model (TFIM), $H(h)=-J\sum_{i=1}^{N-1}\sigma_{i}^{x}\sigma_{i+1}^{x}-h\sum_{i=1}^{N}\sigma_{i}^{z}$, where the interaction strength $J=1$ and the Hamiltonian is parametrized as a function of transverse field, $h$. The inverse temperature $\beta$ is fixed as $1$, the number of ancillas is considered the same as the number of system qubits, and the initial learning rate is taken as $\eta=0.01$ for Meta-VQT and $0.001$ for NN Meta-VQT. The first set of experiments are run on three different system sizes $n=2,3~{}\text{and}~{}4$. The set of Hamiltonian training parameters $h_{train}$ considered are $10$ uniformly spaced points between $-2$ and $2$. The number of encoding and processing layers is chosen as $[2,2],~{}[3,3],~{}\text{and}~{}[4,4]$ for system sizes $n=2,3~{}\text{and}~{}4$, respectively. For all system sizes in NN Meta-VQT, the ansatz configuration has a feed forward neural network consisting of one input, one hidden and two output layers. The output layers from these neural networks provide angles for SU2 hardware efficient and HVA layers. For both Meta and NN Meta-VQT, the test set of Hamiltonian parameters, $h_{test}$ is considered as $40$ uniformly spaced points between $-2$ and $2$. This set also has points that are not present in the $h_{train}$.
Figure 6(a), shows the Gibbs free energy evaluated from the approximate Gibbs state measured from the trained Meta-VQT ansatz for different system sizes. The open circle corresponds to the variational Gibbs free energy evaluated at $h_{train}$ while the cross marker represents the one evaluated for $h_{test}$. The true Gibbs state evaluated analytically in the interval $[-2,2]$ is indicated as dashed lines. The relative error between true Gibbs free energy and the one evaluated from the approximate Gibbs state for the test parameter set is shown in Fig. 6(b). From Figs. 6 (a) and (b), we can conclude that the trained Meta-VQT can generate the Gibbs free energy profile with small relative error for all system sizes for the test set with just $10$ training points. Since the system size considered is relatively smaller, the quality of the prepared Gibbs state is further evaluated by analyzing its fidelity with the exact one as shown in Fig. 6(c). The fidelity stays in the range between $0.93$ to $1$, with a lower fidelity for larger system size.
Figure 7 shows the same set of numerics run with the NN Meta-VQT ansatz. From Fig. 7 (a),(b) and (c), similar to Meta-VQT, NN Meta-VQT also accurately generates the Gibbs free energy profile and Gibbs state for parameter values outside the training set. The relative error in Gibbs free energy (fidelity) for test set is found to be slightly less (more) for states prepared using NN Meta-VQT than Meta-VQT.
The performance of these meta-algorithms as we scale the system size from $2$ to $8$ qubits is shown in Fig. 8(a) and (b). Figure 8(a) plots the mean value with standard deviation of the relative error between the Gibbs free energy evaluated from the trained meta- algorithms and the exact one for values of $h$ in $h_{test}$ as system size increases from $2$ to $8$ qubits. Figure 8(b) shows the mean value with standard deviation of fidelity evaluated between the prepared and true Gibbs state as a function of system size for values of $h$ in $h_{test}$. As the system size increases the relative error (fidelity) increases (decreases). This may be because the $[n,n]$ configuration of encoding (SU2 hardware-efficient) and processing (HVA) layers in Meta-VQT (NN Meta-VQT) not being optimal for all system sizes, requirement of more complex encoding schemes for Hamiltonian parameters and hyperparameter fine tuning such as adjusting the learning rate during the training. Alternatively, a more complex ansatz, such as dynamical PQCs [36], may be required to achieve higher accuracy in generating the Gibbs state for larger system sizes. However, the performance of NN Meta-VQT is slightly better than that of Meta-VQT suggesting that, for larger system sizes, NN Meta-VQT can be scaled more efficiently than Meta-VQT.
Though the performance tends to decrease at larger qubit sizes ($n\geq 6$), our trained meta-algorithms serve as effective initializations for VQT, particularly when optimizing the Gibbs state with respect to a single Hamiltonian parameter. To demonstrate this, we employ a VQT approach using our HEA+HVA ansatz interleaved with CNOT gates, initialized with parameters obtained via meta-training, and optimize the Gibbs free energy corresponding to values of the transverse field parameter $h\in[0,2]$. We consider system sizes of $n=6$ and $n=8$ qubits. For comparison, we evaluate two alternative approaches: (i) the HEA ansatz with a linear CNOT layout and a single ancilla, as proposed in Wang et al.[35], using Gibbs free energy as the cost function; and (ii) the VQT approach used in Verdon et al.[29], Consiglio et al.[30], and related works, which combines two PQCs, for instance an HEA followed by an HVA circuit. In the latter method, intermediate measurements from the first circuit are used to compute the Von Neumann entropy, while the second circuit evaluates the Hamiltonian expectation value. Both quantities are used to construct the Gibbs free energy, and no ancillas are employed. For the meta-trained VQT and the HEA+HVA ansatz without CNOTs in (ii), we use $[n,n]$ layers. For the HEA+CNOT method in (i), we use $n$ layers of HEA. Figure 9(a) and (b) show the fidelity between the exact and variationally prepared Gibbs states for $n=6$ and $n=8$ qubits using three different methods. In both cases, the meta-trained initialization leads to significantly higher fidelity compared to existing VQT approaches, highlighting its advantage in capturing thermal states efficiently for larger systems.
A comparative performance analysis between the proposed meta-algorithms and the existing VarQITE-based Gibbs state preparation is presented in Appendix C, using a 2-qubit Heisenberg model with all field terms included. Even for this small-scale Hamiltonian, the meta-algorithms consistently outperform the VarQITE method. Furthermore, Appendix D provides numerical evidence that the precision of the meta-algorithms remains robust regardless of the number of non-commuting terms in the Hamiltonian, in contrast to VarQITE, whose precision deteriorates as the number of non-commuting terms increases.
Figure 9: VQT initialization with meta-trained ansatz: Fidelity between the exact and variationally prepared Gibbs states for system sizes (a) $n=6$ and (b) $n=8$ qubits, using different implementations of VQT. The meta-trained HEA+HVA ansatz with CNOTs (cyan color) consistently outperforms other approaches, demonstrating the benefit of meta-learning-based initialization in achieving higher fidelity Gibbs state preparation.
Figure 10: Finite-temperature phase diagram of Kitaev ring model : (a) is the magnetic susceptibility, $\chi$, plotted as a function of temperature, $T$, for different h values. For a particular value of $h$, the temperature at which magnetic susceptibility takes maximum value gives the crossover temperature and thus the phase diagram, which is shown in (b). The calculations in both (a) and (b) correspond to system size in the thermodynamic limit. (c) plots the crossover temperature evaluated for finite system sizes $L=2,10,20,50$ along with the one in thermodynamic limit, $L=\infty$. From (c), it is evident that smaller system sizes also gives the phase diagram close to that of the one at $L=\infty$. Figure (a) and (b) are adapted from Shi et al.[40] and (c) is adapted from Zhang et al.[45] .
Figure 11: Preparation of Gibbs state in different regimes of Kitaev ring model : (a) shows the plots of Gibbs free energy evaluated from the approximate Gibbs state prepared using the trained Meta-VQT and VarQITE for different $h$ values of Kitaev ring Hamiltonian. For Meta-VQT the $h$ values considered are $40$ uniform points between $[0.7,1.2]$ including $20$ training points, while for VarQITE it is $20$ uniform points between $[0.7,1.2]$. The black dashed lines correspond to exact Gibbs free energy. The trained Meta-VQT generates Gibbs free energy profile more accurately than VarQITE even for $h$ values which are outside the training set. Additionally, (b) shows the fidelity calculated between the exact Gibbs state and the one approximated using both Meta-VQT and VarQITE. The fidelity for Meta-VQT remains close to one while that of VarQITE depends on the parameter values as well as the regimes of the Hamiltonian. The system size considered is $N=3$. The number of encoding and processing layers for Meta-VQT is $4$ SU2 hardware efficient and $1$ HVA respectively. The depth of SU2 hardware efficient layers in VarQITE is considered to be $5$.
IV.2 Preparing Gibbs state across different regimes of Kitaev ring Hamiltonian
To establish our second claim which states that Meta-VQT and NN Meta-VQT can efficiently simulate the quantum Gibbs state independent of different regimes of the many-body Hamiltonian, we consider a Kitaev ring (a spinless p-wave superconductor) Hamiltonian,
$H(h)=-J\sum_{i=1}^{L-1}\sigma_{i}^{x}\sigma_{i+1}^{x}-J\sigma_{1}^{y}P\sigma_{N}^{y}-h\sum_{i=1}^{L}\sigma_{i}^{z},$ (4)
where, $J=1$, $P=\Pi_{i=2}^{L-1}\sigma_{i}^{z}$ is a string operator and $L$ is the system size. At zero-temperature, the model exhibits a topological quantum phase transition at $h/J=1$ [55].
For finite-temperatures, the Hamiltonian exhibits a temperature crossover behavior as shown in Fig. 10(b). This is characterized by a smooth and gradual transition between semi-classical and quantum critical regimes as a function of temperature and the Hamiltonian’s parameter values [45]. The finite-temperature phase diagram in Fig. 10(b) that shows this crossover behavior is obtained by evaluating the temperature at which magnetic susceptibility, $\chi$, takes maximum value, see Fig. 10(a). Zhang et al.[45] showed that the model exhibits finite-temperature crossover close in range with the infinite system size phase diagram for a wide range of finite system sizes starting from $L=2$, see Fig. 10(c).
Figure 12: Overview of conventional and meta-algorithm-based QBM training : (a) Conventional QBM training: To model a target data distribution $\rho_{data}$, the Gibbs state of a model Hamiltonian $H(J,h)$ must be prepared within each iteration of the QBM training loop. This requires a nested optimization loop where circuit parameters $\theta$ are updated separately for each hamiltonian parameter set $(J,h)$. Once the Gibbs state $\tilde{\rho}(\theta)$ is prepared, the QBM loss $L_{QBM}(\rho_{data},\tilde{\rho}(\theta))$ is minimized by updating $(J,h)$. The dotted pink box highlights the expensive inner-loop Gibbs training step where $\theta$ is optimized. (b) QBM training with meta-algorithm: Here, the variational ansatz is designed to encode the Hamiltonian parameters $(J,h)$ directly to the circuit. The ansatz is initialized with pretrained $\phi_{optimized}$ and $\theta_{optimized}$ obtained from the meat-training stage (as described in Fig. 1(b)). As a result, Gibbs state preparation no longer requires a separate training loop for each $(J,h)$ and the optimization is performed only over $(J,h)$. This eliminates the need for the inner-loop $\theta$ training shown in the dotted pink box in (a), significantly reducing the QPU calls.
Here, we consider a system size of $L=3$ at temperature of $T=1/\beta=0.1$. The Meta-VQT with $6$ total qubits, $4$ encoding and $1$ processing layer is trained for the subsequent generation of quantum Gibbs state as a function of the strength of transverse field $h$. A total of $20$ $h$ values are chosen uniformly from the interval $[0.7,1.2]$ for training purposes. The reason for this particular choice of interval stems from the fact that as $h$ varies from $0.7$ to $1.2$ the model exhibits transition from semi-classical to quantum critical to semi-classical regime for a temperature value of $T=0.1$, see Fig. 10(b) and (c). For testing the Meta-VQT algorithm, $40$ uniform $h$ values are taken from the interval between $[0.7,1.2]$.
The blue open circles in Fig. 11(a) shows the Gibbs free energy evaluated from the Gibbs state prepared using trained Meta-VQT as a function of $h~{}\in~{}[0.7,1.2]$. A comparison of this has been made with the Gibbs free energy evaluated exactly (black dashed lines) and that evaluated from the Gibbs state prepared using VarQITE (red triangle) that has $5$ layers of SU2 hardware efficient ansatz. The Gibbs free energy profile generated from the trained Meta-VQT is more accurate to the exact one than the one evaluated from the states prepared using VarQITE. Moreover, from Fig. 11(b), that plots fidelity between the exact Gibbs states and the ones prepared using trained Meta-VQT (blue open circles) and VarQITE (red triangles) we observe that the Meta-VQT is able to prepare Gibbs state with good fidelity in all finite-temperature regimes of the model while that of VarQITE is dependent on the parameter values as well as the finite-temperature regimes of the model.
IV.3 Application of NN Meta-VQT: Quantum Boltzman machine
In this section, we show one application of our finite-temperature collective optimization algorithm by implementing the NN Meta-VQT in variational QBM workflow.
A QBM relies on preparing finite-temperature quantum Gibbs states of chosen Hamiltonians as a core subroutine. Samples measured from these states are then used to model the target data. An overview on classical and quantum Boltzmann machines are provided in Appendix B. In the traditional variational QBM training workflow (see Fig. 12(a)), once the Hamiltonian and its parameters are initialized, the Gibbs state is prepared using variational methods such as the VQT [29] or VarQITE [33, 11], which employ hybrid quantum-classical optimization. Samples from the prepared Gibbs state are used to evaluate the QBM loss function, typically based on Kullback-Leibler (KL) divergence [10] or quantum relative entropy [12, 13]. If $n_{\text{qbm}}$ are the total epochs required for QBM training and $n_{\text{gibbs}}$ is the average number of optimization steps required for thermal state preparation, then the minimum number of QPU calls required for this algorithm is $n_{\text{qbm}}\times n_{\text{gibbs}}$, assuming that the cost function is evaluated on classical computer and ignoring the shot noise. This implies an additional nested loop for Gibbs state preparation within the QBM training as shown inside pink dotted box in Fig. 12(a). Since our protocol enables collective optimization over a parameterized Hamiltonian, it allows the QBM to bypass this nested loop as illustrated in Fig. 12(b). Consequently, as observed in our numerical examples, the total number of QPU calls reduces to approximately $n_{\text{qbm}}+n_{\text{gibbs}}$ from $n_{\text{qbm}}\times n_{\text{gibbs}}$.
To demonstrate this numerically, we consider generative learning mode of the QBM where the classical probability distribution of a target dataset is modeled by the probability distribution of the Gibbs state of the Hamiltonian. NN Meta-VQT is used for the Gibbs state preparation. The pseudo algorithm for implementing the proposed workflow is shown in Algorithm 3.
The target probability distribution to be learned by the variational QBM is chosen to be $p_{target}=[0.62,0.17,0.17,0.04]$, which corresponds to a random $2$-qubit quantum state. The model Hamiltonian is considered as the $2$-qubit double parameterized Heisenberg Hamiltonian with all field terms in Eq. 7. Since NN Meta-VQT with $4$ processing layers performs the best for this Hamiltonian as observed from Table 3, we use the trained ansatz of NN Meta-VQT of this Hamiltonian for the preparation of the Gibbs state.
Figure 13(a) and (b) shows that during training the loss function of QBM (blue open circles) converges to a lower value, $0.0012$ when NN Meta-VQT is used for preparing Gibbs state as compared to VarQITE based QBM which converges to a higher value of $0.0067$.
Figure 13: NN Meta-VQT in QBM workflow : (a) and (b) shows the convergence of loss function and the trace distance between exact and prepared Gibbs state using our method of the NN Meta-VQT and existing method of the VarQITE respectively. The KL divergence in (a) converges to $0.0012$ while that in (b) converges to $0.0067$. Additionally, the prepared Gibbs state is found to be more accurate and precise for NN Meta-VQT based QBM than VarQITE based QBM. The number of layers considered for NN Meta-VQT is $4$ SU2 while VarQITE requires $32$ layers of SU2 circuit to prepare Gibbs state with an average value of trace distance $0.22\pm 0.094$.
[…]
Although our numerics are based on a 2-qubit Hamiltonian, the timescales indicate that for larger qubit sizes and more complex Hamiltonians and QBMs, the requirement of VarQITE with the SU2 hardware efficient ansatz may necessitate even deeper quantum circuits, potentially increasing the runtime. However, our algorithm, given a smaller timescale, can potentially be scaled to larger system sizes with significantly fewer resources than VarQITE.
V Conclusions
In this work, we introduced two methods of collective optimization known as Meta-VQT and NN Meta-VQT to prepare thermal states of a parametrized Hamiltonian. Unlike the traditional VQT, where the optimization is done for a single parameter configuration of the Hamiltonian, the Meta and NN Meta-VQT is trained to learn the Gibbs state of a set of parameters of the Hamiltonian. This enables efficient generalization across a distribution of Hamiltonians, reducing quantum resource overhead, and allowing rapid inference of thermal states for unseen configurations without retraining from scratch. We accomplish this by encoding the parameters of the Hamiltonian using an encoding layer, which is a SU2 hardware efficient ansatz in Meta-VQT and a neural network in NN Meta-VQT, followed by a processing layer which is HVA in Meta-VQT (Fig. 4) and hardware efficient ansatz and///or HVA in NN Meta-VQT (Fig. 5). The trainable angles in Meta-VQT is provided by the encoding and processing quantum circuits, while for NN Meta-VQT the trainable parameters are in the classical neural network. We observe that the ansatz (Meta-VQT and NN Meta-VQT) trained for a set of $10$ parameters of the Hamiltonian is able to generate Gibbs state close to the exact one corresponding to all other parameters not in the training set, but within the range of parameters chosen for training.
We demonstrate this using a parametrized TFIM Hamiltonian in Fig. 6 and Fig. 7. For larger system sizes, our meta-trained ansatz can be optimized for Gibbs state corresponding to single parameter of the Hamiltonian and outperforms two other existing randomly initialized variational techniques as shown in Fig. 9. The example of a $2$-qubit double parametrized Heisenberg Hamiltonian with external field shows that our method of collective optimization is able to prepare Gibbs state with better precision than the existing technique of VarQITE, as tabulated in Table 3. Additionally, unlike VarQITE, the precision and accuracy of prepared Gibbs state from Meta and NN Meta-VQT remains independent of the complexity of the Hamiltonian as shown in Table  2. Furthermore, because of the unique structure of the ansatz chosen for Meta and NN Meta-VQT, the Algorithm 1 and Algorithm  2 is used to simulate quantum systems that exhibits finite-temperature crossover. This is demonstrated using a Kitaev ring model that exhibits temperature crossover close to the exact one for small system sizes such as $L=3$. The existing literature has separate ansatzes and algorithms for simulating the different regimes across the crossover temperature such as semi-classical and quantum critical regimes with quantum critical being the most difficult to simulate. However, here we show numerically that our unique choice of ansatz is able to learn the features of Gibbs state of all the regimes of the Hamiltonian. When compared to VarQITE, the accuracy and precision of Gibbs state prepared using Meta-VQT remains better in all parameter regimes of the Kitaev ring model, see Fig. 11. Additionally, collective optimization techniques are known to overcome some of the bottlenecks in VQE [42] such as barren plateaus [56, 57, 58] and local minima [59, 60] or non-convexity in the loss function landscape, and we expect that the Meta-VQT and NN Meta-VQT can also achieve this. However, further work needs to be done to confirm this and is beyond the scope of this paper.
Finally, we demonstrate an important application of our new algorithm for Gibbs state preparation in the variational QBM workflow. Our meta-algorithm based QBM presents several important advantages over traditional VQT-based QBM training. Firstly, the nested loop required otherwise for preparing Gibbs state of the Hamiltonian in variational QBM workflow is avoided. The meta-trained ansatz is used to prepare the Gibbs state for all parameters of Hamiltonian in QBM training. This reduces the quantum resources required and runtime of the QBM training. Secondly, since the Meta and NN Meta-VQT algorithm can be trained independent of the different regimes of the Hamiltonian, complex many-body Hamiltonians can be used to model the target distribution without affecting the quality of prepared Gibbs state and hence the overall performance of QBM. In the numerical experiments, using NN Meta-VQT in QBM, we are able to bring down the runtime by upto $30$ times as compared to the existing technique of training QBM using VarQITE. Also, the Gibbs state prepared during the entire training process is more precise and the loss function of the QBM converges to a lower value for NN Meta-VQT based QBM, see Fig. 13.
VI Open questions
There are some challenges that must be addressed while implementing the proposed Meta and NN Meta-VQT algorithms in this work. The major one arises when the size of the Hamiltonian whose Gibbs state needs to be prepared increases beyond $N\geq 10$. In this case, the evaluation of exact Gibbs free energy becomes challenging due to the exponential increase in size of the density matrix that represents the Gibbs state and its memory requirement. Alternate methods can be used to evaluate the Gibbs free energy as the one explained in Verdon et al.[29] or Wang et al. [35]. Furthermore, the number of encoding and processing layers required by Meta-VQT increases with the system size, which places an onus on the need to have a detailed study on trainability of these models in light of scalability and barren plateau analysis, and for potential improvements for these algorithms in the future.
Some possible approaches for mitigating these issues include developing a booster method for Meta and NN Meta-VQT similar to the ones designed for VQE by Wang et al.[61], designing the ansatz with barren plateau free PQC’s as mentioned in Deshpande et al.[36], incorporating multi-qubit non-unitary operations devised by Za- pusek et al.[37] as well as utilizing training schemes that reduces resource costs on quantum hardware as introduced by Bhowmick et al.[62]. The neural network chosen for NN Meta-VQT also provides a channel for the scope of improvement when dealing with more complex and larger size Hamiltonians.
Another fundamental challenge, involves the theoretical hardness of Gibbs state preparation itself. For general Hamiltonians, preparing the Gibbs state is a QMA-hard problem. Thus, identifying classes of Hamiltonians for which efficient variational Gibbs state preparation is feasible remains an open theoretical and practical challenge. Relatedly, for applications like QBMs, the problem of selecting a suitable Hamiltonian that can faithfully and efficiently capture the underlying data distribution is still not well understood and continues to be an active area of research.
In this work, our comparative study for Gibbs state preparation has been done with only few existing variational technique in literature for specific depth and $\beta$ values. Though the numerical experiments suggest that the performance of Meta VQT and NN Meta-VQT is better than that of VarQITE under the specific conditions chosen for our numerical study, there is still scope for comparison with other methods.
The extension of these studies in the continuous variable domain is a possible future direction to explore. A further important consideration is conducting error analysis and resource estimation for scaled versions of these models, particularly in the context of the noisy hardware on which they are trained or implemented. Additionally, a comprehensive understanding of the boundaries of quantum utility or advantage for such heuristic methods remains an open question, requiring further theoretical advancements.
VII Acknowledgement
The authors thank Dr. Quoc Hoan Tran from the Quantum Laboratory at Fujitsu Research Japan, Dr. Bibhas Adhikari and Dr. Hannes Leipold from Quantum Laboratory at Fujitsu Research America for their valuable feedback at different stages of the work and for taking the time to review the current draft. The authors extend their immense gratitude to Yasuhiro Endo, Hirotaka Oshima and Shintaro Sato as well as the entire Robust Quantum Computing Department at Fujitsu Limited for their strategic and technical support. The code used in this paper will be made available at a later date. A preliminary version of this paper was presented at the Fujitsu-IISc Quantum Workshop in Bangalore, India, and a poster presentation is scheduled for TQC 2025, to be held from September 15–19.
Appendix A Variational Quantum Imaginary Time Evolution (VarQITE)
In this method, the approximate Gibbs state is evaluated using PQCs that implements variational quantum imaginary time evolution based on McLachlan’s variational principle [33, 11]. A brief explanation of how VarQITE is used to approximate the Gibbs state of the Hamiltonian is presented below.
For an initial state, |$\ket{\psi_{0}}$, the imaginary time $(\tau=it)$ evolution (ITE) with respect to the Hamiltonian, $H(\vec{h})$, non-unitarily maps the state to,
|$\ket{\psi_{\tau}}=\frac{e^{-H(\vec{h})\tau}\ket{\psi_{0}}}{\sqrt{\bra{\psi_{0}}e^{-2H(\vec{h})\tau}\ket{\psi_{0}}}},$ (5)
and follows Wick-rotated Schrodinger’s equation,
$\frac{d\ket{\psi_{\tau}}}{d\tau}=-(H(\vec{h})-E_{\tau})\ket{\psi_{\tau}},$
where $E_{\tau}=\bra{\psi_{\tau}}H(\vec{h})\ket{\psi_{\tau}}$ is the system’s energy. If the initial state |$\ket{\psi_{0}}$ is taken to be a purification of the maximally mixed state, $\rho=\text{Tr}_{A}\ket{\psi_{0}}\bra{\psi_{0}}=\mathbb{I}/d$, where A represents the ancilla subsystem, and the imaginary time evolution is kept as $\tau\rightarrow 1/2T=\beta/2$ for the system, then Eq. (5) yields a purification of the Gibbs state defined in Eq. (1).
The ITE state cannot be directly implemented on a quantum circuit because of the non-unitary nature of evolution. However, McArdle et al.[33] showed that the ITE can be implemented on a hybrid gate-based quantum computer using McLachlan’s variational principle. The main idea here is as follows. First, the state |$\ket{\psi_{\tau}}$ is approximated by defining a parametrized trial state : |$\ket{\phi(\vec{\omega}(\tau))}=U(\vec{\omega}(\tau))\ket{0}$, where $U(\vec{\omega}(\tau))=U_{N}(\omega_{n})U_{N-1}(\omega_{N-1}).....U_{k}(\omega_{k})....U_{1}(\omega_{1})$ are the set of parametrized single or two-qubit gates with parameters $\vec{\omega}(\tau)$. Since the goal is the preparation of Gibbs state and this requires purification of maximally mixed states as the initial boundary condition, the state |$\ket{\phi(\vec{\omega}(\tau))}$ is defined over a PQC with a total of $\text{n}+\text{n}_{\text{ancilla}}$ qubits. The initial value of parameters of the unitaries are chosen such that,
|$\ket{\phi(\vec{\omega}(\tau=0))}=U(\vec{\omega}(\tau=0))\ket{0}^{{\otimes}2n}=\ket{\phi^{+}}^{\otimes n},$
where |$\ket{\phi^{+}}^{\otimes n}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})$ is the Bell state. Here, in each Bell state |$\ket{\phi^{+}}$, the first qubit belongs to the system, while the second qubit serves as an ancilla. Tracing out the ancillas from |$\ket{\phi(\omega(0))}$ results in the maximally mixed state; $\frac{\mathbb{I}}{2^{n}}$. Further, the ITE of |$\ket{\phi(\vec{\omega}(\tau))}$ is done by implementing McLachlan variational principle,
$\delta||(\frac{\partial}{\partial\tau}+H(\vec{h})-E_{\tau})\ket{\phi({\vec{\omega}(\tau)})}||=0,$
using quantum circuits for $\tau=\beta/2$. The state, $\text{Tr}_{b}[\ket{\phi(\vec{\omega}(\tau=\beta/2))}]$ at the end of $\tau=\beta/2$ evolution gives the approximation to the Gibbs state of our system Hamiltonian, $H(\vec{h})$. For more detailed calculations, please refer to Zoufal et al. [11] and McArdle et al. [33]. In this work, the technique has been adapted solely for comparative study using specific examples, with the variational quantum imaginary time evolution simulated using the built-in library available in Qiskit.
Appendix B Overview of Classical and Quantum Boltzmann Machines
This section provides brief overview of classical and quantum Boltzmann machines, both of which fundamentally rely on the system’s Gibbs state.
B.1 Classical Boltzmann Machine
Classical Boltzmann Machines are a type of energy-based stochastic neural network inspired by statistical mechanics. They learn complex probability distributions over input data. The network consists of interconnected nodes of visible and hidden neurons, that commonly represent binary states. The input data is represented using the visible neurons while the latent features in the input data is learned by the hidden neurons. Figure 14, shows a type of Boltzmann machine known as Restricted Boltzmann Machine (RBM) where connections exist between visible and hidden neurons but not among neurons within the same layer.
Figure 14: Restricted Boltzmann Machine : The figure illustrates the architecture of a classical restricted Boltzmann machine, where connections exist between visible and hidden neurons but not among neurons within the same layer. $a_{i}$ and $b_{i}$ are the biases of visible and hidden neurons, and $w_{i,j}$ are the connection weights between hidden and visible neurons. The input data is mapped to the visible neurons. The hidden neurons learn the latent features in the input data by adjusting the connection weights and biases. The output is then obtained by sampling the visible layer after a round-trip of Gibbs sampling, where information is propagated from the visible to hidden layer and then back to the visible layer. This reconstructed visible state serves as the model’s output, enabling it to generate samples, reconstruct inputs, or estimate probability distributions over input data.
The entire network is described by an energy function, $E(v,h)=-\sum_{i=1}^{N_{v}}a_{i}v_{i}-\sum_{i=1}^{N_{h}}b_{i}h_{i}-\sum_{i,j}w_{i,j}v_{i}h_{j}$, where $v$ is the visible unit, $h$ is the hidden unit, $N_{v}$ represents the number of visible units, $N_{h}$ represents the number of hidden units, $a_{i}$ and $b_{i}$ are the visible and hidden unit biases, and $w_{i,j}$ is the connection weight between visible and hidden neurons. The probability distribution over possible unit configurations is represented by the Boltzmann distribution, $P(v,h)=e^{-\beta E(v,h)}/\sum_{v,h}e^{-\beta E(v,h)}$, which characterizes the state of the system at finite temperature, $T$. During training, the temperature controls the level of stochasticity in the model, with higher temperatures leading to more uniform sampling across configurations, while lower temperatures bias the system toward lower energy (more probable) states. The weights connecting the visible and hidden neurons and the biases are trained so that the probability distribution of the target input data is accurately represented by $P(v)=\sum_{h}e^{-\beta E(v,h)}/\sum_{v,h}e^{-\beta E(v,h)}$. Boltzmann machines are used in various applications, including dimensionality reduction [63], feature learning [64], collaborative filtering [65], and pretraining of deep learning networks [66].
B.2 Quantum Boltzmann Machine (QBM)
The quantum version of this classical architecture is known as the Quantum Boltzmann Machine. The energy function representing the network structure of classical BM shown in Fig. 14, is replaced by the parametrized Hamiltonian,
$H(\vec{h})=\sum_{i=0}^{p-1}h_{i}\sigma_{i},$ (6)
where visible and hidden units are now qubits defined by the Pauli matrices $\sigma^{x},\sigma^{y}$ and $\sigma^{z}$ , $\vec{h}=(h^{1},h^{2},...h^{p})\,\epsilon\,\mathbb{R}^{p}$ are the trainable coefficients of the Hamiltonian and $\sigma_{i}=\otimes_{j=0}^{n-1}\sigma_{i}^{j}$ with $\sigma$’s as the Pauli matrices and $\sigma_{i}^{j}\,\epsilon\,{I,X,Y,Z}$. The visible qubits are those that are used to model the input data and measure the model output, while the hidden qubits are those that represent the latent features. In analogy with classical BM, the aim of QBM is to learn the parameters, $\vec{h}$, such that the probability distribution of the Gibbs state of the Hamiltonian $H(\vec{h})$, approximates the target distributions. With the inclusion of non-commuting terms in the Hamiltonian which are absent in classical Boltzmann machines, QBMs have the potential for greater expressive power, enabling them to model a broader class of complex data distributions.
For a configuration |$\ket{v}$ of the visible qubits, a projective measurement on this visible state is represented as $\hat{\Pi}_{\mathbf{v}}=\ket{v}\bra{v}^{\mathbf{v}}\otimes\mathbb{I}^{\mathbf{h}}$, where $\mathbf{v}$ and $\mathbf{h}$ indicate the visible and hidden subsystems, respectively. The marginal probability of observing the visible configuration $\mathbf{v}$ is then given by,
$p_{\mathbf{v}}(\vec{h})=\text{Tr}[\hat{\Pi}_{\mathbf{v}}\rho(\vec{h})],$
where $\rho(\vec{h})=e^{-\beta\hat{H}(\vec{h})}/\text{Tr}[e^{-\beta\hat{H}(\vec{h})}]$ is the Gibbs state of the full Hamiltonian which includes both the visible and hidden qubits. If $p^{\text{data}}$ is the probability distribution that represents the input data, then the training of QBM involves minimizing the distance between $p^{\text{data}}$ and $p_{\mathbf{v}}(\vec{h})$. The distance between two probability distributions is typically quantified by the Kullback-Leibler (KL) divergence,
$\mathcal{L}(\vec{h})=\sum_{\mathbf{v}}p^{\text{data}}\log\frac{p^{\text{data}}}{p_{\mathbf{v}}(\vec{h})}.$
The KL divergence always takes a nonnegative value, $L(\vec{h})\geq 0$, with $L(\vec{h})=0$ when $p^{\text{data}}=p_{\mathbf{v}}(\vec{h})$. When the input data is provided as an encoded quantum state $\rho^{data}$, the relative quantum entropy $S(\rho^{\text{data}}||\rho(\vec{h}))=\text{Tr}[\rho^{\text{data}}\log\rho^{\text{data}}-\rho^{\text{data}}\log\rho(\vec{h})]$, is used to quantify the difference between the QBM-modeled data and the input target data [3]. Minimization of the KL divergence or relative entropy is done on a classical [41] or quantum hardware [11] by evaluating their respective first-order derivatives with respect to coefficient $\vec{h}$ of the Hamiltonian [10]. Furthermore, the optimization is carried out by updating the Hamiltonian parameters based on these gradients as follows,
$\vec{h}^{\prime}=\vec{h}+\eta\partial_{\vec{h}}\mathcal{L}(\vec{h}),$
where $\eta$ is the learning rate.
In this paper, we consider training the QBM with only visible units using our proposed meta-algorithms for Gibbs state preparation. The computation of the gradients of the loss function is done using numerical differentiation.
Table 2: Robustness to Hamiltonian complexity : We compare the quality of Gibbs states prepared by NN Meta-VQT and VarQITE as a function of the complexity of a $2$-qubit Hamiltonian. The precision of NN Meta-VQT, measured by the avearge trace distance between the prepared and true Gibbs state over a range of $(J,h)$ values, remains largely unaffected as the number of non-commuting terms increases, demonstrating robustness to Hamiltonian complexity. In contrast, the accuracy of VarQITE degrades noticeably when the Hamiltonian contains more than two non-commuting operator blocks.
Hamiltonian Commuting Blocks NN Meta-VQT VarQITE
$H=-J\sigma_{1}^{x}\sigma_{2}^{x}$ $1$ $\mathbf{0.04\pm 0.01}$ $0.002\pm 0.001$
$H=-J\sigma_{1}^{x}\sigma_{2}^{x}-h\sum_{i=1}^{2}\sigma_{i}^{x}$ 1 $\mathbf{0.05\pm 0.02}$ $0.005\pm 0.003$
$H=-J\sigma_{1}^{x}\sigma_{2}^{x}-h\sum_{i=1}^{2}(\sigma_{i}^{x}+\sigma_{i}^{z})$ 2 $\mathbf{0.06\pm 0.02}$ $0.016\pm 0.006$
$H=-J\sigma_{1}^{x}\sigma_{2}^{x}-h\sum_{i=1}^{2}(\sigma_{i}^{x}+\sigma_{i}^{z}+\sigma_{i}^{y})$ 3 $\mathbf{0.07\pm 0.05}$ $0.33\pm 0.20$
$H=-J(\sigma_{1}^{x}\sigma_{2}^{x}+\sigma_{1}^{y}\sigma_{2}^{y})-h\sum_{i=1}^{2}(\sigma_{i}^{x}+\sigma_{i}^{z}+\sigma_{i}^{y})$ 3 $\mathbf{0.08\pm 0.05}$ $0.30\pm 0.17$
$H=-J(\sigma_{1}^{x}\sigma_{2}^{x}+\sigma_{1}^{y}\sigma_{2}^{y}+\sigma_{1}^{z}\sigma_{2}^{z})-h\sum_{i=1}^{2}(\sigma_{i}^{x}+\sigma_{i}^{z}+\sigma_{i}^{y})$ 3 $\mathbf{0.08\pm 0.03}$ $0.38\pm 0.17$
Table 3: Comparsion between Meta-VQT, NN Meta-VQT and VarQITE : For the 2-qubit Heisenberg Hamiltonian with external fields shown in Eq. (7), the Meta-VQT and NN Meta-VQT algorithms prepare the Gibbs state with much better precision than VarQITE technique with similar circuit depth.
Method Layers Average Trace distance
Meta-VQT 4 SU2 + 1 HVA $0.18\pm 0.07$
Meta-VQT 4 SU2 $0.22\pm 0.09$
NN Meta-VQT 4 SU2 + 1 HVA $0.15\pm 0.02$
NN Meta-VQT 4SU2 $0.08\pm 0.03$
VarQITE 4SU2 $0.32\pm 0.20$
Appendix C Preparing Gibbs state of a Heisenberg Hamiltonian with external field
To further expand the scope, the Meta and NN Meta- VQT is trained to generate the quantum Gibbs state of a $2$-qubit Heisenberg Hamiltonian with external fields,
$H(J,h)=-J(\sigma_{1}^{x}\sigma_{2}^{x}+\sigma_{1}^{y}\sigma_{2}^{y}+\sigma_{1}^{z}\sigma_{2}^{z})-\\ h(\sum_{i=1}^{2}(\sigma_{i}^{z}+\sigma_{i}^{x}+\sigma_{i}^{y})).$ (7)
The parameter values used for training Meta-VQT (NN Meta-VQT) are $10$ uniformly (randomly) selected values between $-2$ and $2$ for each parameter $J$ and $h$. The average trace distance between the prepared and exact Gibbs state is evaluated and compared with the VarQITE technique existing in literature as shown in Table 3.
The quantum circuit depth chosen for VarQITE is $4$ layers of SU2 hardware efficient ansatz. For Meta-VQT and NN Meta-VQT, two cases are considered, one where the ansatz is a $4$ layered SU2 hardware efficient circuit and the other with $4$ SU2 and $1$ HVA layers. The number of ancillas required by the two meta-algorithms and VarQITE is same as the system size which is $2$. From Table 3, the average trace distance between the true Gibbs state and the one evaluated from Meta and NN Meta-VQT algorithm for all four cases is lower and more precise than the avearge trace distance evaluated between true Gibbs state and the one prepared using VarQITE. Additionally, the NN Meta-VQT algorithm generates good quality Gibbs state even in the absence of HVA, while the Meta-VQT with at least one additional layer of HVA can improve the quality of Gibbs state than the one without HVA.
Appendix D Complexity of the Hamiltonian and the efficiency of NN Meta-VQT
In this section, we present results from another numerical study that does an analysis of the dependency of quality of prepared Gibbs state using different algorithms and the complexity of the Hamiltonian. For the numerics, the NN Meta-VQT with $4$ SU2 hardware efficient layers is considered for an equivalent comparison with the VarQITE that also have $4$ SU2 hardware efficient layers. The complexity of the Hamiltonian is analyzed using number of commuting blocks, which also indicates the number of non-commuting terms in the Hamiltonian [67]. The average trace distance between the target and prepared Gibbs state from VarQITE and NN Meta-VQT as a function of complexity of Hamiltonian is shown in Table 2. The efficiency measured as the average trace distance between true and prepared Gibbs state of VarQITE decreases as the Hamiltonian becomes more complex while for the NN Meta-VQT the efficiency remains independent of the complexity of the Hamiltonian. Additionally, for Hamiltonian with more than $2$ commuting blocks the NN Meta-VQT efficiency is found to be better than that of the VarQITE.
Appendix E Pseudoalgorithms
The pseudoalgorithms to run Meta-VQT, NN Meta-VQT and NN Meta-VQT based QBM are shown in Algorithm 1, Algorithm 2 and Algorithm 3, respectively.
Algorithm 1 Meta-Variational Quantum Thermalizer
Input: Parameter $\vec{h}$ of Hamiltonian
Output: Quantum Gibbs state of $H(\vec{h})$
// Initialization
$\beta\leftarrow 1$ // Inverse temperature
$n_{\text{ancilla}}\leftarrow n_{\text{system}}$ // Number of ancillas
$n_{\text{proc}}\leftarrow n_{\text{enco}}\leftarrow n_{\text{system}}$ // Number of circuit layers
$\vec{h}_{\text{train}}$ // Training parameter set
$n_{\text{epochs}}$ // Number of training steps
$\eta\leftarrow 0.01$ // Learning rate
$(\vec{\omega},\vec{\theta},\vec{\phi})\leftarrow(\vec{\omega}_{0},\vec{\theta}_{0},\vec{\phi}_{0})$ // Initial circuit parameters
// Training Loop
for $n=0$ to $n_{\text{epochs}}$ do
     Loss $\leftarrow 0$
     for all $\vec{h}_{i}\in\vec{h}_{\text{train}}$ do
[…]
         ProcAnsatz $\leftarrow\mathrm{HVA}(\vec{\theta})$
         Ansatz $\leftarrow$ EncoAnsatz + ProcAnsatz
[…]
         $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
         $G(\vec{h}_{i})\leftarrow\text{Evaluate Gibbs free energy}(\tilde{\rho})$
         Loss $\leftarrow$ Loss + $G(\vec{h}_{i})$
     end for
     Global loss $\leftarrow$ Loss
[…]
     $(\vec{\omega},\vec{\theta},\vec{\phi})\leftarrow(\vec{\omega},\vec{\theta},\vec{\phi})-\eta\cdot\text{Gradients}$
end for
return $\text{Ansatz}(\vec{\omega},\vec{\theta},\vec{\phi})$
Algorithm 2 Neural Network Meta-Variational Quantum Thermalizer
Input: Parameter $\vec{h}$ of Hamiltonian
Output: Quantum Gibbs state of $H(\vec{h})$
// Initialization
$\beta\leftarrow 1$ // Inverse temperature
$n_{\text{ancilla}}\leftarrow n_{\text{system}}$ // Number of ancillas
$n_{\text{proc}}\leftarrow n_{\text{enco}}\leftarrow n_{\text{system}}$  // Number of circuit layers
$\vec{h}_{\text{train}}$ // Training parameter set
$n_{\text{epochs}}$ // Number of training steps
$\eta\leftarrow 0.001$ // Learning rate
$\vec{\phi}_{\text{nn}}\leftarrow\vec{\phi}^{0}_{\text{nn}}$ // Initial neural network parameters
// Training Loop
for $n=0$ to $n_{\text{epochs}}$ do
     $\text{Loss}\leftarrow 0$
     for $\vec{h}_{i}\in\vec{h}_{\text{train}}$ do
         Encoding neural $\text{Encoding neural network}\leftarrow(\vec{h}_{i},\vec{\phi}_{nn})$
         $\vec{\theta}\leftarrow\text{Encoding neural network}(\vec{h}_{i},\vec{\phi}_{\text{nn}})$
[…]
         $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
         $G(\vec{h}_{i})\leftarrow\text{Evaluate Gibbs free energy}(\tilde{\rho})$
         $\text{Loss}\leftarrow\text{Loss}+G(\vec{h}_{i})$
     end for
[…]
     $\vec{\phi}_{\text{nn}}\leftarrow\vec{\phi}_{\text{nn}}-\eta\cdot\text{Gradients}$
end for
// Return final circuit
return $\text{Ansatz}(\vec{\phi}_{\text{nn}})$
Algorithm 3 Variational QBM with NN Meta-VQT
Input: Probability distribution of a random quantum state
Output: Hamiltonian whose Gibbs states models the input distribution.
// Initialization
$p_{target}$ // Choose target probability distribution
$H(\vec{J},\vec{h})$ // Choose model Hamiltonian
$(\vec{J},\vec{h})$ // Trainable coefficients of Hamiltonian
$\text{Ansatz}(\vec{\phi}_{nn})$ // Trained NN Meta-VQT ansatz from Algorithm 2
$n_{epochs}$ // Number of training steps
$\eta\leftarrow 0.1$ // Learning rate
// Training Loop
for $n=0$ to $n_{epochs}$ do
     Ansatz $(\vec{\phi}_{nn})$leftarrow $(\vec{h},\vec{J})$
     Output state $\leftarrow$ Ansatz $(\vec{\phi}_{nn})(\vec{h},\vec{J})$
[…]
     $p_{var}\leftarrow$ Measure diagonal elements$(\tilde{\rho})$
     Loss $\leftarrow$ KL divergence $(p_{var},p_{target})$
     Gradients $\leftarrow$ Compute gradients (Loss)
     $(\vec{J},\vec{h})\leftarrow(\vec{J},\vec{h})-\eta\cdot\text{Gradients}$.
end for
return $(\vec{J},\vec{h})$

### v2

@@
 I Introduction
-Generation of finite-temperature thermal states or Gibbs states of a given Hamiltonian is an integral task in state preparation using quantum computers.
-Gibbs states find applications in a variety of tasks such as quantum simulation [2], quantum machine learning [3, 4], quantum optimization [5], and the study of open quantum systems [6].
-In particular, combinatorial optimization problems [5], semi-definite programming [7], quantum key distribution [8], quantum metrology [9], and training of Quantum Boltzmann Machines (QBMs) [3, 10, 11, 12, 13, 14, 15] are tackled by sampling from well-engineered Gibbs states.
-Preparing arbitrary quantum states is generally a challenging task; in particular, finding the ground state of a Hamiltonian is known to be QMA-hard [16].
-Preparing Gibbs states, especially at low and intermediate temperatures, is as hard as finding the ground state of that Hamiltonian [17].
-Various methods propose preparing Gibbs states using classical and quantum computers [18, 6, 19, 20, 21, 22, 23, 24].
-Some proposals that focus on quantum computers include quantum rejection sampling [25], quantum walk [26], and dynamics simulation [27, 28].
-Although these methods may incur exponential cost in the worst case, they are efficient in special scenarios such as when the ratio between the partition functions of the infinite temperature state and the Gibbs state is at most polynomially large [6], and when the gap of the Markov chain describing the quantum walk is polynomially small [27, 26].
-However, these algorithms suit fault-tolerant quantum computers as they require complex quantum subroutines like quantum phase estimation.
+Preparation of quantum Gibbs states is of fundamental importance as they provide statistical description of quantum systems at finite temperature, which is the natural setting for almost all real-world physical systems.
+Beyond their foundational role in statistical mechanics such as the study of open quantum systems [2], Gibbs states have emerged as valuable resources in quantum technologies, underpinning applications in areas such as quantum simulation [3], quantum machine learning [4, 5, 6, 7, 8, 9, 10, 11], optimization [12], quantum metrology [13] and quantum cryptography [14].
+Their broad utility, coupled with the intrinsic difficulty of preparing them efficiently [15], has motivated an extensive line of research aimed at developing algorithms and protocols for Gibbs state generation using classical and quantum computational techniques [16, 2, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
 Figure 1: Overview of meta-algorithm for Gibbs state preparation :(a) Depicts the core task addressed by our meta-algorithm: preparing thermal states for a parametrized many-body spin Hamiltonian at finite temperature.
@@
 These are then used to reconstruct the Gibbs free energy profile.
+While many of these existing quantum computing approaches are designed with fault-tolerant quantum hardware in mind, near-term devices are constrained by noise and limited circuit depth.
 Variational Quantum Algorithms (VQAs) that can be implemented efficiently on noisy intermediate-scale quantum (NISQ) devices, present an alternative to these methods.
-VQAs are a class of hybrid quantum-classical algorithms that optimize a loss function depending on the circuit parameters.
-The loss function is evaluated on quantum hardware, while parameter updates are computed on classical processors.
+VQAs represent a class of hybrid algorithms that combine quantum and classical resources.
+The quantum device evaluates a loss function determined by the circuit parameters, and a classical optimizer updates these parameters iteratively.
 Several variational techniques have also been proposed for Gibbs state preparation on near-term quantum hardware.
@@
 Qubits
-Quantum Hamiltonian-Based Models and the Variational Quantum Thermalizer Algorithm, Verdon et al. [29] $4$ 0
-Variational quantum Gibbs state preparation with a truncated Taylor series, Wang et al.[35] 11 1
-Extending the Variational Quantum Eigensolver to Finite Temperatures, Selisko et al.[31] 4 0
-Variational quantum simulation of the quantum critical regime, Shi et al. [40] 6 0
-Training Quantum Boltzmann Machines with the $\beta$ Variational Quantum Eigensolver, Huijgen et al.[41] 10 0
-Thermal state preparation of the SYK model using a variational quantum algorithm, Araz et al.[32] 12 0
-Dynamic parameterized quantum circuits: expressive and barren-plateau free, Deshpande et al.[36] 10 0
-Meta and Neural Network Meta-Variational Quantum Thermalizer, Our Work 8 8
+Quantum Hamiltonian-Based Models and the Variational Quantum Thermalizer Algorithm,
+Verdon et al. [27]
+44 0
+Variational quantum Gibbs state preparation with a truncated Taylor series,
+Wang et al.[33]
+11 1
+Extending the Variational Quantum Eigensolver to Finite Temperatures,
+Selisko et al.[29]
+4 0
+Variational quantum simulation of the quantum critical regime,
+Shi et al. [38]
+6 0
+Training Quantum Boltzmann Machines with the $\beta$ Variational Quantum Eigensolver,
+Huijgen et al.[39]
+10 0
+Thermal state preparation of the SYK model using a variational quantum algorithm,
+Araz et al.[30]
+12 0
+Dynamic parameterized quantum circuits: expressive and barren-plateau free,
+Deshpande et al.[34]
+10 0
+Meta and Neural Network Meta-Variational Quantum Thermalizer,
+Our Work
+8 8
 In this work, we address these challenges by proposing two novel algorithms :
@@
 For convenience, we assume that the Boltzmann constant is equal to $1$ throughout the paper.
-Given a PQC represented by unitaries $U(\vec{\theta})$ with $m$ trainable angles, $\vec{\theta}\in\theta_{1},\theta_{2},...\theta_{m}$ acting on $\text{n}+\text{n}_{\text{ancilla}}$ qubits, the goal of VQT is to train the PQC to approximate the Gibbs state of this Hamiltonian as, $\rho(\vec{h})\approx\tilde{\rho}(\vec{\theta})=\text{Tr}_{\text{ancilla}}\left(\ket{\psi(\vec{\theta})}\bra{\psi(\vec{\theta})}\right)$, where |$\ket{\psi(\vec{\theta})}=U(\vec{\theta})\ket{\psi_{0}}$ is the state generated by the PQC from the initial state |$\ket{\psi_{0}}$.
-The exact loss function to be minimized using a classical computer to train this PQC is the Gibbs free energy given by,
+[…]
 $G(\vec{\theta},\vec{h})=\langle H(\vec{h})\rangle-TS,$ (2)
@@
 The minimization of $G(\vec{\theta},\vec{h})$ ensures that the system reaches the corresponding Gibbs state which is the state of minimal energy at finite-temperature.
+The schematic workflow of VQT is shown in Fig.
+2(a).
 This is in analogy with the VQE where the expectation value of the Hamiltonian is minimized to find the ground state energy.
@@
 However, unlike VQE, the mixed nature of the Gibbs state necessitates the use of ancilla qubits $(1~\leq~\text{n}_{\text{ancilla}}~\leq~\text{n})$ to prepare the approximate Gibbs state of given Hamiltonian.
-The required ancilla qubits can scale up further depending on how the loss function is evaluated on a quantum computer.
-For instance, Wang et al.[35] employs a truncated Taylor series expansion of order $3$ that needs a SWAP test to measure the approximate Von-Neumann entropy, thus requiring multiple copies of the quantum state and consequentially more ancillas.
-The schematic workflow of VQT is shown in Fig.
-2(a).
+II.1.1 Challenges in Gibbs Free Energy Evaluation and Scalable Approaches
+Evaluating the exact Gibbs free energy becomes highly challenging for large system sizes.
+The main difficulty arises from the entropy contribution in Eq.
+2, whose computation is generally intractable even on quantum computers for arbitrary large quantum states[45, 46].
+To address this, several works have proposed algorithms that approximate the Gibbs free energy while maintaining scalability on near-term devices.
+Two notable approaches are:
+  • The work by Wang et al.[33] expands the entropy in a Taylor series and truncates it at some finite order $K=2$ or $K=3$.
+The truncated series of entropy corresponds to evaluating higher-order overlaps between quantum states.
+Consequently, Gibbs free energy can be estimated as a linear combination of the system’s energy and these state overlaps, which are efficiently measurable on near-term quantum hardware.
+The trade-off is that evaluating higher-order state overlaps requires additional ancilla qubits.
+  • Johannes Selisko et al.[29] introduced the quantum VQT (qVQT) algorithm, which employs two variational quantum circuits (VQCs) together with an intermediate measurement to prepare a mixed state that approximates the Gibbs state.
+The first VQC produces a probability distribution via intermediate measurement, enabling entropy estimation, while the second VQC evaluates the system’s energy.
+A classical optimizer iteratively tunes both parameter sets to minimize the Gibbs free energy.
+This hybrid quantum-classical scheme provides a resource-efficient and scalable route to Gibbs state preparation, adaptable to different ansatz families.
+Related approaches have also been applied by Verdon et al.
+[27] and J.
+Y.
+Araz [30] to prepare Gibbs states for models such as the 2D Heisenberg Hamiltonian and the Sachdev–Ye–Kitaev (SYK) model.
 Figure 4: Meta-VQT ansatz : The Hamiltonian parameters, $\vec{h}$, are encoded in the angles of the single qubit parametrized gates through a linear transformation.
@@
 It finds applications in tasks such as generalization of ground states of parametrized Hamiltonian (Meta-VQE and Neural Network (NN) Meta-VQE or quantum circuit learning) [1, 40, 41, 42], training algorithms that does not rely on gradient computation [50] and training classical neural networks to assist in the quantum learning process [51, 52].
-In this work, we introduce the finite-temperature versions of Meta-VQE and NN Meta-VQE, known as Meta-VQT and NN Meta-VQT, which generates the Gibbs free energy profile, $G(\vec{h})$, and prepares the corresponding quantum Gibbs state, $\rho(\vec{h})$, of a parametrized Hamiltonian, $H(\vec{h})$.
-The workflow of our proposed Meta-VQT/NN Meta-VQT is shown in Fig.
-2(b).
-The finite-temperature meta-algorithms are developed as generalized versions of the VQT algorithm explained in Sec.
-II.1.
-Here, the PQC, |$\ket{\psi(\vec{\theta},\vec{\phi},\vec{\omega})}$ (or neural network) is trained for multiple discrete input parameters ${\vec{h}_{train}}\in[\vec{h}_{1},\vec{h}_{2},...,\vec{h}_{n}]$ of the Hamiltonian, $H(\vec{h})$, so that for the optimal ansatz parameters, $(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})$ (or $\vec{\phi}_{nn}^{*})$, the quantum circuit is able to generate quantum Gibbs state for Hamiltonian parameters $\vec{h}_{test}\notin{\vec{h}_{train}}$.
-Suppose that $\vec{h}$ is a single-valued vector $h$ and $h_{train}\in(h_{1},h_{2},.....h_{n})$ is the training set with $h$ values arranged in ascending order.
-Then the meta-algorithms perform well for $h_{test}\notin h_{train}$ and $h_{1}\leq h_{test}\leq h_{n}$, whereas outside this range, its performance tends to decline.
-The number of distinctive trainable angles depends on the type of meta-algorithm, size of $\vec{h}$, the type of encoding and the class of PQC used for training.
+[…]
 Since, the target Gibbs state is usually mixed, additional ancilla qubits are required to prepare quantum Gibbs state that we trace out before measurement.
@@
 However, for simpler Hamiltonians, fewer ancilla qubits also demonstrates high accuracy.
-Figure 5: NN Meta-VQT ansatz : In NN Meta-VQT, the neural network encodes the Hamiltonian parameters, $\vec{h}$, and provides the circuit parameters, $\vec{\theta}$, as output.
-The quantum circuit is constructed out of both hardware-efficient and HVA as shown above.
-Similar to Meta-VQT, after this final processing layer, measurement on the system qubits yield a density matrix, $\tilde{\rho}(\vec{\theta})$, that is further optimized to approximate the true Gibbs state of the system.
-The loss function used is the sum of Gibbs free energy corresponding to each parameter in $\vec{h}_{train}$.
-For system sizes considered here, where the number of qubits ranges from $n=2~{}\text{to}~{}8$, the Gibbs free energy is evaluated on a classical computer after measuring the Gibbs state, $\rho(\vec{h})\approx\tilde{\rho}(\vec{\theta})=\text{Tr}_{\text{ancilla}}\left(\ket{\psi(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})}\bra{\psi(\vec{\theta}^{*},\vec{\phi}^{*},\vec{\omega}^{*})}\right)$, from the quantum circuit.
-For larger system sizes, where the size of the Gibbs state becomes too large to be stored in a classical computer, the Gibbs free energy itself is evaluated from the quantum circuits using a truncated Taylor series expansion of the entropy [35] or by applying the method described by Verdon et al.
-[29].
+[…]
 The following sections outline the details regarding the Meta-VQT and NN Meta-VQT algorithms.
@@
 The number of trainable parameters, $\vec{\theta}$, in HVA depends on the terms in the Hamiltonian whose Gibbs state needs to be prepared.
-Figure 6: Gibbs state preparation of TFIM using Meta-VQT : The results of Meta-VQT algorithm run for different qubit sizes of TFIM is shown.
-The value of $\beta$ is chosen to be $1$.
-(a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set.
-The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of Meta-VQT is shown in (b).
-The relative error for all qubit sizes are low indicating the high efficiency of trained Meta-VQT in predicting the Gibbs free energy for test parameters outside the training set.
-This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from the trained Meta-VQT for test parameters gives value between $0.93$ to $1$.
-The number of encoding and processing layers chosen for system sizes $n=2,3,4$ are $[2,2],[3,3]~{}\text{and}~{}[4,4]$ respectively.
+Figure 5: NN Meta-VQT ansatz : In NN Meta-VQT, the neural network encodes the Hamiltonian parameters, $\vec{h}$, and provides the circuit parameters, $\vec{\theta}$, as output.
+The quantum circuit is constructed out of both hardware-efficient and HVA as shown above.
+Similar to Meta-VQT, after this final processing layer, measurement on the system qubits yield a density matrix, $\tilde{\rho}(\vec{\theta})$, that is further optimized to approximate the true Gibbs state of the system.
 The density matrix, $\tilde{\rho}(\vec{\omega},\vec{\phi},\vec{\theta},\vec{h})$, measured by tracing out the ancillas is used to evaluate the Gibbs free energy, $G(\vec{h})$, of $H(\vec{h})$.
@@
 III.1.2 Neural Network Meta-Variational Quantum Thermalizer (NN Meta-VQT)
+Figure 6: Gibbs state preparation of TFIM using Meta-VQT : The results of Meta-VQT algorithm run for different qubit sizes of TFIM is shown.
+The value of $\beta$ is chosen to be 11.
+(a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set.
+The relative error between exact Gibbs free energy and the Gibbs free energy evaluated using variationally prepared Gibbs state at $h_{test}$ with trained parameters of Meta-VQT is shown in (b).
+The relative error for all qubit sizes are low indicating the high efficiency of trained Meta-VQT in predicting the Gibbs free energy for test parameters outside the training set.
+This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from the trained Meta-VQT for test parameters gives value between $0.93$ to 11.
+The number of encoding and processing layers chosen for system sizes $n=2,3,4$ are $2,2],[3,3]~\text{and}~[4,4]$ respectively.
 Neural Network Meta-VQT (NN Meta-VQT) builds on the previous variant by adding a classical neural network to the quantum circuit as shown in Step 3 (green colored box) of Fig.
@@
 Figure 7: Gibbs state preparation of TFIM using NN Meta-VQT : The results of NN Meta-VQT algorithm for different qubit sizes of TFIM is shown.
-The value of $\beta$ is chosen to be $1$.
+The value of $\beta$ is chosen to be 11.
 (a) shows the exact Gibbs free energy (dashed) evaluated in the interval $h\in[-2,2]$, and the one calculated after training from the variationally prepared Gibbs state for the parameters which are present (open circles) and absent (cross mark) in the training set.
@@
 The relative error for all qubit sizes are low indicating the high efficiency of trained neural network architecture in predicting the Gibbs free energy for test parameters outside the training set.
-This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from trained NN Meta-VQT ansatz for test parameters gives value between $0.98$ to $1$ for $n=2,3$ and $4$.
-The number of SU2 hardware efficient and HVA ansatz layers in the processing part is chosen as $[2,2],~{}[3,3]$ and $[4,4]$ for system sizes $n=2,3$ and $4$ respectively.
+This is also evident from (c) where the fidelity between exact Gibbs state and Gibbs state measured from trained NN Meta-VQT ansatz for test parameters gives value between $0.98$ to 11 for $n=2,3$ and $4$.
+The number of SU2 hardware efficient and HVA ansatz layers in the processing part is chosen as [2,2],[3,3][2,2],~[3,3] and $[4,4]$ for system sizes $n=2,3$ and 44 respectively.
 Figure 8: Scaling of Meta-VQT and NN Meta-VQT:(a) shows the mean value with standard deviation of relative error between the Gibbs free energy evaluated for values of $h$ in $h_{test}$ from the trained meta-algorithms and the exact one for system sizes $2,4,6$ and $8$ qubits.
@@
 For $n$ qubit Hamiltonian, the number of encoding (SU2 hardware efficient) and processing (HVA) layers considered for Meta-VQT (NN Meta-VQT) is $[n,n]$.
-The test set, $h_{test}$, over which the mean and standard deviation in relative error in Gibbs free energy and fidelity is evaluated are $40$ uniform points between $[-2,2]$.
+The test set, $h_{test}$, over which the mean and standard deviation in relative error in Gibbs free energy and fidelity is evaluated are $40$ uniform points between [2,2][-2,2].
 IV Numerical Results
@@
 The first set of experiments are run on three different system sizes $n=2,3~\text{and}~4$.
-The set of Hamiltonian training parameters $h_{train}$ considered are $10$ uniformly spaced points between $-2$ and $2$.
-The number of encoding and processing layers is chosen as $[2,2],~{}[3,3],~{}\text{and}~{}[4,4]$ for system sizes $n=2,3~{}\text{and}~{}4$, respectively.
+The set of Hamiltonian training parameters $h_{train}$ considered are $10$ uniformly spaced points between 2-2 and $2$.
+The number of encoding and processing layers is chosen as $2,2],~[3,3],~\text{and}~[4,4]$ for system sizes $n=2,3~\text{and}~4$, respectively.
 For all system sizes in NN Meta-VQT, the ansatz configuration has a feed forward neural network consisting of one input, one hidden and two output layers.
 The output layers from these neural networks provide angles for SU2 hardware efficient and HVA layers.
-For both Meta and NN Meta-VQT, the test set of Hamiltonian parameters, $h_{test}$ is considered as $40$ uniformly spaced points between $-2$ and $2$.
+For both Meta and NN Meta-VQT, the test set of Hamiltonian parameters, $h_{test}$ is considered as $40$ uniformly spaced points between 2-2 and $2$.
 This set also has points that are not present in the $h_{train}$.
@@
 The open circle corresponds to the variational Gibbs free energy evaluated at $h_{train}$ while the cross marker represents the one evaluated for $h_{test}$.
-The true Gibbs state evaluated analytically in the interval $[-2,2]$ is indicated as dashed lines.
+The true Gibbs state evaluated analytically in the interval [2,2][-2,2] is indicated as dashed lines.
 The relative error between true Gibbs free energy and the one evaluated from the approximate Gibbs state for the test parameter set is shown in Fig.
@@
 Figure 11: Preparation of Gibbs state in different regimes of Kitaev ring model : (a) shows the plots of Gibbs free energy evaluated from the approximate Gibbs state prepared using the trained Meta-VQT and VarQITE for different $h$ values of Kitaev ring Hamiltonian.
-For Meta-VQT the $h$ values considered are $40$ uniform points between $[0.7,1.2]$ including $20$ training points, while for VarQITE it is $20$ uniform points between $[0.7,1.2]$.
+For Meta-VQT the $h$ values considered are $40$ uniform points between $[0.7,1.2]$ including 2020 training points, while for VarQITE it is 2020 uniform points between $[0.7,1.2]$.
 The black dashed lines correspond to exact Gibbs free energy.
@@
 The system size considered is $N=3$.
-The number of encoding and processing layers for Meta-VQT is $4$ SU2 hardware efficient and $1$ HVA respectively.
-The depth of SU2 hardware efficient layers in VarQITE is considered to be $5$.
+The number of encoding and processing layers for Meta-VQT is 44 SU2 hardware efficient and $1$ HVA respectively.
+The depth of SU2 hardware efficient layers in VarQITE is considered to be 55.
 IV.2 Preparing Gibbs state across different regimes of Kitaev ring Hamiltonian
@@
 The Meta-VQT with $6$ total qubits, $4$ encoding and $1$ processing layer is trained for the subsequent generation of quantum Gibbs state as a function of the strength of transverse field $h$.
-A total of $20$ $h$ values are chosen uniformly from the interval $[0.7,1.2]$ for training purposes.
+A total of 2020 $h$ values are chosen uniformly from the interval $[0.7,1.2]$ for training purposes.
 The reason for this particular choice of interval stems from the fact that as $h$ varies from $0.7$ to $1.2$ the model exhibits transition from semi-classical to quantum critical to semi-classical regime for a temperature value of $T=0.1$, see Fig.
@@
 Additionally, the prepared Gibbs state is found to be more accurate and precise for NN Meta-VQT based QBM than VarQITE based QBM.
-The number of layers considered for NN Meta-VQT is $4$ SU2 while VarQITE requires $32$ layers of SU2 circuit to prepare Gibbs state with an average value of trace distance $0.22\pm 0.094$.
-[…]
+The number of layers considered for NN Meta-VQT is 44 SU2 while VarQITE requires $32$ layers of SU2 circuit to prepare Gibbs state with an average value of trace distance $0.22\pm 0.094$.
+During the training phase, the trace distance between the exact and prepared Gibbs state from NN Meta-VQT and VarQITE (red triangles) is plotted in Fig.
+13(a) and (b).
+For NN Meta-VQT based QBM, the average trace distance during the training phase is $0.071\pm 0.013$, and remains much lower than the one for VarQITE based QBM that has a value of $0.22\pm 0.094$.
+The circuit depth for VarQITE is $32$, which is also much larger than the depth required by NN Meta-VQT to achieve this precision.
+The runtime for training NN Meta-VQT is $\approx 40~\text{minutes}$ and with this trained ansatz the QBM training takes $16~\text{seconds}$.
+However, for VarQITE based QBM the total runtime is around $1200~\text{minutes}$ which is $30$ times more than the total runtime of NN Meta-VQT based QBM which takes $\approx 40~\text{minutes}~16~\text{seconds}$.
 Although our numerics are based on a 2-qubit Hamiltonian, the timescales indicate that for larger qubit sizes and more complex Hamiltonians and QBMs, the requirement of VarQITE with the SU2 hardware efficient ansatz may necessitate even deeper quantum circuits, potentially increasing the runtime.
@@
 We accomplish this by encoding the parameters of the Hamiltonian using an encoding layer, which is a SU2 hardware efficient ansatz in Meta-VQT and a neural network in NN Meta-VQT, followed by a processing layer which is HVA in Meta-VQT (Fig.
-4) and hardware efficient ansatz and///or HVA in NN Meta-VQT (Fig.
+4) and hardware efficient ansatz and//or HVA in NN Meta-VQT (Fig.
 5).
@@
 A brief explanation of how VarQITE is used to approximate the Gibbs state of the Hamiltonian is presented below.
-For an initial state, |$\ket{\psi_{0}}$, the imaginary time $(\tau=it)$ evolution (ITE) with respect to the Hamiltonian, $H(\vec{h})$, non-unitarily maps the state to,
-|$\ket{\psi_{\tau}}=\frac{e^{-H(\vec{h})\tau}\ket{\psi_{0}}}{\sqrt{\bra{\psi_{0}}e^{-2H(\vec{h})\tau}\ket{\psi_{0}}}},$ (5)
+[…]
+[…]
 and follows Wick-rotated Schrodinger’s equation,
-$\frac{d\ket{\psi_{\tau}}}{d\tau}=-(H(\vec{h})-E_{\tau})\ket{\psi_{\tau}},$
-where $E_{\tau}=\bra{\psi_{\tau}}H(\vec{h})\ket{\psi_{\tau}}$ is the system’s energy.
-If the initial state |$\ket{\psi_{0}}$ is taken to be a purification of the maximally mixed state, $\rho=\text{Tr}_{A}\ket{\psi_{0}}\bra{\psi_{0}}=\mathbb{I}/d$, where A represents the ancilla subsystem, and the imaginary time evolution is kept as $\tau\rightarrow 1/2T=\beta/2$ for the system, then Eq.
-(5) yields a purification of the Gibbs state defined in Eq.
-(1).
-The ITE state cannot be directly implemented on a quantum circuit because of the non-unitary nature of evolution.
-However, McArdle et al.[33] showed that the ITE can be implemented on a hybrid gate-based quantum computer using McLachlan’s variational principle.
-The main idea here is as follows.
-First, the state |$\ket{\psi_{\tau}}$ is approximated by defining a parametrized trial state : |$\ket{\phi(\vec{\omega}(\tau))}=U(\vec{\omega}(\tau))\ket{0}$, where $U(\vec{\omega}(\tau))=U_{N}(\omega_{n})U_{N-1}(\omega_{N-1}).....U_{k}(\omega_{k})....U_{1}(\omega_{1})$ are the set of parametrized single or two-qubit gates with parameters $\vec{\omega}(\tau)$.
-Since the goal is the preparation of Gibbs state and this requires purification of maximally mixed states as the initial boundary condition, the state |$\ket{\phi(\vec{\omega}(\tau))}$ is defined over a PQC with a total of $\text{n}+\text{n}_{\text{ancilla}}$ qubits.
-The initial value of parameters of the unitaries are chosen such that,
-|$\ket{\phi(\vec{\omega}(\tau=0))}=U(\vec{\omega}(\tau=0))\ket{0}^{{\otimes}2n}=\ket{\phi^{+}}^{\otimes n},$
-where |$\ket{\phi^{+}}^{\otimes n}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})$ is the Bell state.
-Here, in each Bell state |$\ket{\phi^{+}}$, the first qubit belongs to the system, while the second qubit serves as an ancilla.
-Tracing out the ancillas from |$\ket{\phi(\omega(0))}$ results in the maximally mixed state; $\frac{\mathbb{I}}{2^{n}}$.
-Further, the ITE of |$\ket{\phi(\vec{\omega}(\tau))}$ is done by implementing McLachlan variational principle,
-$\delta||(\frac{\partial}{\partial\tau}+H(\vec{h})-E_{\tau})\ket{\phi({\vec{\omega}(\tau)})}||=0,$
-using quantum circuits for $\tau=\beta/2$.
-The state, $\text{Tr}_{b}[\ket{\phi(\vec{\omega}(\tau=\beta/2))}]$ at the end of $\tau=\beta/2$ evolution gives the approximation to the Gibbs state of our system Hamiltonian, $H(\vec{h})$.
-For more detailed calculations, please refer to Zoufal et al.
-[11] and McArdle et al.
-[33].
-In this work, the technique has been adapted solely for comparative study using specific examples, with the variational quantum imaginary time evolution simulated using the built-in library available in Qiskit.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Appendix B Overview of Classical and Quantum Boltzmann Machines
@@
 With the inclusion of non-commuting terms in the Hamiltonian which are absent in classical Boltzmann machines, QBMs have the potential for greater expressive power, enabling them to model a broader class of complex data distributions.
-For a configuration |$\ket{v}$ of the visible qubits, a projective measurement on this visible state is represented as $\hat{\Pi}_{\mathbf{v}}=\ket{v}\bra{v}^{\mathbf{v}}\otimes\mathbb{I}^{\mathbf{h}}$, where $\mathbf{v}$ and $\mathbf{h}$ indicate the visible and hidden subsystems, respectively.
-The marginal probability of observing the visible configuration $\mathbf{v}$ is then given by,
+[…]
 $p_{\mathbf{v}}(\vec{h})=\text{Tr}[\hat{\Pi}_{\mathbf{v}}\rho(\vec{h})],$
@@
 Hamiltonian Commuting Blocks NN Meta-VQT VarQITE
-$H=-J\sigma_{1}^{x}\sigma_{2}^{x}$ $1$ $\mathbf{0.04\pm 0.01}$ $0.002\pm 0.001$
+$H=-J\sigma_{1}^{x}\sigma_{2}^{x}$ 11 $\mathbf{0.04\pm 0.01}$ $0.002\pm 0.001$
 $H=-J\sigma_{1}^{x}\sigma_{2}^{x}-h\sum_{i=1}^{2}\sigma_{i}^{x}$ 1 $\mathbf{0.05\pm 0.02}$ $0.005\pm 0.003$
@@
 $H(J,h)=-J(\sigma_{1}^{x}\sigma_{2}^{x}+\sigma_{1}^{y}\sigma_{2}^{y}+\sigma_{1}^{z}\sigma_{2}^{z})-\\ h(\sum_{i=1}^{2}(\sigma_{i}^{z}+\sigma_{i}^{x}+\sigma_{i}^{y})).$ (7)
-The parameter values used for training Meta-VQT (NN Meta-VQT) are $10$ uniformly (randomly) selected values between $-2$ and $2$ for each parameter $J$ and $h$.
+The parameter values used for training Meta-VQT (NN Meta-VQT) are $10$ uniformly (randomly) selected values between 2-2 and $2$ for each parameter $J$ and $h$.
 The average trace distance between the prepared and exact Gibbs state is evaluated and compared with the VarQITE technique existing in literature as shown in Table 3.
@@
   for all $\vec{h}_{i}\in\vec{h}_{\text{train}}$ do
-[…]
+   EncoAnsatz $\leftarrow\mathrm{SU2}(\vec{h}_{i},\vec{\omega},\vec{\phi})$
    ProcAnsatz $\leftarrow\mathrm{HVA}(\vec{\theta})$
    Ansatz $\leftarrow$ EncoAnsatz + ProcAnsatz
+   Output state $\leftarrow\text{Simulate}(\text{Ansatz})$
+   $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
 […]
-         $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
-         $G(\vec{h}_{i})\leftarrow\text{Evaluate Gibbs free energy}(\tilde{\rho})$
    Loss $\leftarrow$ Loss + $G(\vec{h}_{i})$
@@
   Global loss $\leftarrow$ Loss
-[…]
+  Gradients $\leftarrow\text{Compute gradients}(\text{Global loss})$
   $(\vec{\omega},\vec{\theta},\vec{\phi})\leftarrow(\vec{\omega},\vec{\theta},\vec{\phi})-\eta\cdot\text{Gradients}$
@@
   for $\vec{h}_{i}\in\vec{h}_{\text{train}}$ do
-         Encoding neural $\text{Encoding neural network}\leftarrow(\vec{h}_{i},\vec{\phi}_{nn})$
+   $\text{Encoding neural network}\leftarrow(\vec{h}_{i},\vec{\phi}_{nn})$
    $\vec{\theta}\leftarrow\text{Encoding neural network}(\vec{h}_{i},\vec{\phi}_{\text{nn}})$
+   $\text{Ansatz}\leftarrow\text{SU2}(\vec{\theta}^{(1)})+\text{HVA}(\vec{\theta}^{(2)})$
+   $\text{Output state}\leftarrow\text{Simulate}(\text{Ansatz})$
+   $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
 […]
-[…]
-         $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
-         $G(\vec{h}_{i})\leftarrow\text{Evaluate Gibbs free energy}(\tilde{\rho})$
    $\text{Loss}\leftarrow\text{Loss}+G(\vec{h}_{i})$
   end for
-[…]
-[…]
+  $\text{Global loss}\leftarrow\text{Loss}$
+  $\text{Gradients}\leftarrow\text{Compute gradients}(\text{Global loss})$
   $\vec{\phi}_{\text{nn}}\leftarrow\vec{\phi}_{\text{nn}}-\eta\cdot\text{Gradients}$
@@
 for $n=0$ to $n_{epochs}$ do
-     Ansatz $(\vec{\phi}_{nn})$leftarrow $(\vec{h},\vec{J})$
+  Ansatz $(\vec{\phi}_{nn})$ $\leftarrow$ $(\vec{h},\vec{J})$
   Output state $\leftarrow$ Ansatz $(\vec{\phi}_{nn})(\vec{h},\vec{J})$
+  $\tilde{\rho}\leftarrow$ Trace out ancilla (Output State)($\text${Output State})
 […]
-     $p_{var}\leftarrow$ Measure diagonal elements$(\tilde{\rho})$
   Loss $\leftarrow$ KL divergence $(p_{var},p_{target})$